### v1

# Maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$

Abstract.
We introduce maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ whose deformation space can be understood as the cotangent bundle of Weil-Petersson universal Teichmüller space $T_{0}(1)$. We show that the Mess map is a symplectic diffeomorphism from $T^{*}T_{0}(1)$ to $T_{0}(1)\times T_{0}(1)$ for a canonical symplectic form on $T^{*}T_{0}(1)$ and the difference of two Weil-Petersson symplectic forms on $T_{0}(1)\times T_{0}(1)$. Moreover, we show that the anti-holomorphic energy functional of the induced Gauss map is a Kähler potential of the canonical symplectic form over certain subspaces of $T^{*}T_{0}(1)$.
2010 Mathematics Subject Classification 30F60, 32G15, 53C50, 53A10 .
1. Introduction
In the seminal work in [15], Takhtajan and Teo introduced a subspace of the universal Teichmüller space $T(1)$, which is equipped with a Hermitian inner product. This space denoted by $T_{0}(1)$ is the Weil-Petersson universal Teichmüller space in the sense that the space $T_{0}(1)$ has rich structures derived from the Weil-Petersson inner product and the associated symplectic 2-form $\omega_{\mathrm{WP}}$. Moreover, in [15], they also introduced the universal Liouville action $S$ for an element in $T_{0}(1)$ and proved that $S$ is a Kähler potential function of $\omega_{\mathrm{WP}}$ such that
(1.1) ¯$\partial\bar{\partial}S=-2\sqrt{-1}\,\omega_{\mathrm{WP}}\qquad\text{over}\quad T_{0}(1).$
Here $\partial$ and $\bar{\partial}$ denote the holomorphic and anti-holomorphic derivatives on $T_{0}(1)$ respectively. The equality (1.1) can be considered as the universal version of the results which have been proved for the classical Liouville actions over Teichmüller spaces of Riemann surfaces in [16], [17], [14], [12].
An element in $T_{0}(1)$ can be represented as a quasi-circle in the complex plane with a certain condition, which is called Weil-Petersson curve. After its introduction in [15], this has drawn many attentions and has become a very active research object in various research subjects. In particular, Wang in [23] introduced the Loewner energy $I_{L}$ for a Weil-Petersson curve using the energy of driving function which naturally appears in the Shramm Loewner equation and showed the following equality
(1.2) $S=\pi I_{L}\qquad\text{over}\quad T_{0}(1).$
This equality is rather surprising since the same quantity can be defined in terms of totally different constructions developed in Teichmüller theory and Schramm Loewner theory.
In a recent work [3] about the Weil-Petersson curve, Bishop extensively studied various aspects of Weil-Petersson curves in the view points of analysis and geometry. In particular, in [3], Bishop considered an invariant for a minimal surface in the hyperbolic 3-space $\mathbb{H}^{3}$ bounding a given Weil-Petersson curve in ¯$\partial\overline{\mathbb{H}}^{3}$. One drawback of this approach is the non-uniqueness of the bounding minimal surfaces in $\mathbb{H}^{3}$. Hence, it seems to be difficult to relate an invariant of a bounding minimal surface to the universal Liouville action or equivalently the Loewner energy of the Weil-Petersson curve.
Motivated by the works in [15], [23], and [3], we consider a maximal surface in the Anti de Sitter 3-space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, which is a Lorentzian counter part of a minimal surface in $\mathbb{H}^{3}$. More precisely, we introduce a maximal disc of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ such that its boundary is the graph of a quasisymmetric homeomorphism of $\mathbb{S}^{1}$ representing an element in $T_{0}(1)$. By the work [4], such a bounding maximal surface exists uniquely for a given element in $T_{0}(1)$. Allowing a deformation of conformal structures in $T_{0}(1)$ on the given (real) maximal disc, we can see that the cotangent space $T^{*}T_{0}(1)$ parametrizes the embedded maximal discs of Weil-Petersson class in the Anti de Sitter 3-space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. Actually, we do not need the Weil-Petersson condition to apply the result of [4]. But, it turns out that the Weil-Petersson condition is very essential for results in this paper, which we are going to explain.
One of the basic properties for a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is the existence of a pair of harmonic maps $F_{\pm}$ induced by the Gauss map of $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. In general, it is not possible to get a finite valued energy of the harmonic maps $F_{\pm}$ since the domain $\Sigma$ is not compact. But, in Proposition 3.7, it is proved that the Weil-Petersson condition for a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ implies the finiteness of the anti-holomorphic energy of the harmonic maps $F_{\pm}$ on $\Sigma$.
Following the original construction of Mess in his work [10], one can introduce so called Mess map
(1.3) $\mathrm{Mess}:T^{*}T(1)\to T(1)\times T(1).$
It has been expected that the map $\mathrm{Mess}$ in (1.3) is a symplectic diffeomorphism. We prove this property for the restriction of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$. Recall that the Weil-Petersson symplectic form $\omega_{\mathrm{WP}}$ is defined on $T_{0}(1)$ by using the Hilbert manifold structure on $T_{0}(1)$ in [15]. It is also crucial to use the Hilbert manifold structure on $T_{0}(1)$ to introduce the canonical symplectic form $\omega_{\mathrm{C}}$ on $T^{*}T_{0}(1)$. Then, our result for this states as follows
Theorem 1.1.
The map $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is a symplectic diffeomorphism, that is,
$\omega_{\mathrm{C}}=-\mathrm{Mess}^{*}_{+}(\omega_{\mathrm{WP}})+\mathrm{Mess}^{*}_{-}(\omega_{\mathrm{WP}}).$
Here $\mathrm{Mess}_{\pm}:=\pi_{\pm}\circ\mathrm{Mess}$ where $\pi_{\pm}$ denotes the projection map from $T_{0}(1)\times T_{0}(1)$ onto the first or second factor respectively.
A related result to Theorem 1.1 was given in [7] at the origin of $T^{*}T(1)$ and $T(1)\times T(1)$. But, it is not clear which symplectic structures are referred to in [7] considering that a main result of [15] is the construction of the symplectic form $\omega_{\mathrm{WP}}$ on $T_{0}(1)$.
Let us put $T_{0}(1)^{+}$, $T_{0}(1)^{-}$ to be the inverse image of $T_{0}(1)\times\{0\}$, $\{0\}\times T_{0}(1)$ by the map $\mathrm{Mess}$ respectively. Then we have
Theorem 1.2.
The anti-holomorphic energy $E(F_{\pm})$ of the induced harmonic maps $F_{\pm}$ defines a finite valued functional over $T^{*}T_{0}(1)$. Moreover, we have
(1.4) $2\,\partial\bar{\partial}E=\mp\sqrt{-1}\,i^{*}_{\pm}\,\omega_{\mathrm{C}}\qquad\text{over}\quad T_{0}(1)^{\pm}.$
Here $i_{\pm}:T_{0}(1)^{\pm}\to T^{*}T_{0}(1)$ denotes the inclusion map of $T_{0}(1)^{\pm}$ respectively.
Combining (1.1), (1.2), Theorem 1.1, and Theorem 1.2, we have
Corollary 1.3.
The following equality holds over $T_{0}(1)^{\pm}$,
(1.5) $4\,\partial\bar{\partial}E=i^{*}_{\pm}\mathrm{Mess}_{\pm}^{*}(\partial\bar{\partial}S)=\pi\,i^{*}_{\pm}\mathrm{Mess}_{\pm}^{*}(\partial\bar{\partial}I_{L}).$
Corollary 1.3 states that the anti-holomorphic energy functional plays the same role as the Liouville action $S$ or equivalently the Loewner energy $\pi I_{L}$ up to the symplectic diffeomorphism $\mathrm{Mess}$.
Here is an explanation of structure of this paper. In Section 2, we introduce some basic materials on the universal Teichmüller space $T(1)$ and the Weil-Petersson universal Teichmüller space $T_{0}(1)$. In Section 3, we introduce the definition of the maximal disc of Weil-Petersson class and study some properties of it. In Section 4 and 5, we prove main theorems using the variational methods which have been used in [16], [17], [14], [15], [12], and [11]. In the appendix, we present some basics about the 3-dimensional Anti de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$.
2. Universal Teichmüller Space
This section is a brief introduction of the universal Teichmüller space and the Weil-Petersson universal Teichmüller space. We refer to the chapter 1 of [9] and the chapter 16 of [6] for the universal Teichmüller space and the chapter 1 of [15] for the Weil-Petersson universal Teichmüller space.
2.1. Universal Teichmüller space
Let $\mathrm{QS}(\mathbb{S}^{1})$ denote the group of the quasisymmetric homeomorphism of the circle $\mathbb{S}^{1}$. Then the universal Teichmüller space is defined by
(2.1) $T(1):=\mathrm{Mob}(\mathbb{S}^{1})\backslash\mathrm{QS}(\mathbb{S}^{1})$
where $\mathrm{Mob}(\mathbb{S}^{1})\cong\mathrm{PSL}(2,\mathbb{R})$ denotes the subgroup preserving $\mathbb{S}^{1}$ of the Möbius transformation group $\mathrm{PSL}(2,\mathbb{C})$ acting on ${}^\widehat{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$.
Let $\mathbb{D}=\{\,z\in\mathbb{C}\,|\,|z|<1\,\}$ be the open unit disk and let $\mathbb{D}^{*}=\{\,z\in\mathbb{C}\,|\,|z|>1\,\}$ be its exterior in $\widehat{\mathbb{C}}$. Denote by $L^{\infty}(\mathbb{D})$ and $L^{\infty}(\mathbb{D}^{*})$ the complex Banach spaces of bounded Beltrami differentials on $\mathbb{D}$ and $\mathbb{D}^{*}$ respectively, and let $L^{\infty}(\mathbb{D})_{1}$ be the unit ball in $L^{\infty}(\mathbb{D})$. Given a Beltrami differential $\mu\in L^{\infty}(\mathbb{D})_{1}$, extend it to $\mathbb{D}^{*}$ by the reflection
(2.2) $\mu(z)=\overline{\mu\Big{(}\frac{1}{\bar{z}}\Big{)}}\frac{z^{2}}{\bar{z}^{2}},\qquad\text{for}\quad z\in\mathbb{D}^{*}.$
Let $w_{\mu}:\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ be the solution of the Beltrami differential equation
(2.3) $\partial_{\bar{z}}w_{\mu}=\mu\,\partial_{z}w_{\mu}$
with fixed points $1,-1,-\sqrt{-1}$. Then $w_{\mu}$ preserves $\mathbb{S}^{1}$ and $w_{\mu}|_{\mathbb{S}^{1}}\in\mathrm{QS}(\mathbb{S}^{1})$. Conversely, by extension theorem of Beurling-Ahlfors, any quasisymmetric homeomorphism in $\mathrm{QS}(\mathbb{S}^{1})$ can be extended to a quasiconformal homeomorphism $w_{\mu}$ of $\mathbb{D}$ for a $\mu\in L^{\infty}(\mathbb{D})_{1}$, hence we have the following description of $T(1)$,
(2.4) $T(1)=L^{\infty}(\mathbb{D})_{1}/\sim$
where $\mu\sim\nu$ if and only if $w_{\mu}|_{\mathbb{S}^{1}}=w_{\nu}|_{\mathbb{S}^{1}}$. We denote the equivalence class of $\mu$ by $[\mu]\in T(1)$. The space $T(1)$ has a unique structure of a complex Banach manifold such that the projection map
$\mathrm{P}:L^{\infty}(\mathbb{D})_{1}\to T(1)$
is a holomorphic submersion. The differential of $\mathrm{P}$ at the origin
$D_{0}\mathrm{P}:L^{\infty}(\mathbb{D})\to T_{[0]}T(1)$
is a complex linear surjection of holomorphic tangent spaces. The kernel of $D_{0}\mathrm{P}$ is the subspace $\mathcal{N}(\mathbb{D})$ of infinitesimally trivial Beltrami differentials.
For a given Beltrami differential $\mu\in L^{\infty}(\mathbb{D})_{1}$, extend it to be zero on $\mathbb{D}^{*}$. Let $w^{\mu}$ be the unique solution to the Beltrami differential equation
$\partial_{\bar{z}}w^{\mu}=\mu\,\partial_{z}w^{\mu}$
with the normalization $w^{\mu}(0)=0$, $(w^{\mu})_{z}(0)=1$ and $(w^{\mu})_{zz}(0)=0$. Then $w^{\mu}$ is conformal on $\mathbb{D}^{*}$. Then, we have
(2.5) $T(1)=L^{\infty}(\mathbb{D})_{1}/\sim$
where $\mu\sim\nu$ if and only if $w^{\mu}|_{\mathbb{D}}=w^{\nu}|_{\mathbb{D}}$. This characterization is equivalent to the one in (2.4) since $w_{\mu}|_{\mathbb{S}^{1}}=w_{\nu}|_{\mathbb{S}^{1}}$ if and only if $w^{\mu}|_{\mathbb{D}}=w^{\nu}|_{\mathbb{D}}$.
Now, we define the Bers embedding of $T(1)$ into the complex Banach space
$A_{\infty}(\mathbb{D}^{*})=\{\,\phi:\mathbb{D}^{*}\to\mathbb{C}\,|\,\text{holomorphic},\ \mathrm{sup}_{\mathbb{D}^{*}}|\phi|e^{-\psi}<\infty\,\}$
where $e^{\psi}$ denotes the hyperbolic density function on $\mathbb{D}^{*}$. For a holomorphic map $f$ on an open domain in $\widehat{\mathbb{C}}$, the Schwarzian of $f$ is defined by
(2.6) $\mathcal{S}(f)=\left(\frac{f_{zz}}{f_{z}}\right)_{z}-\frac{1}{2}\left(\frac{f_{zz}}{f_{z}}\right)^{2}.$
For every $\mu\in L^{\infty}(\mathbb{D})_{1}$, the holomorphic function $\mathcal{S}(w^{\mu}|_{\mathbb{D}^{*}})$ lies in $A_{\infty}(\mathbb{D}^{*})$ by Kraus-Nehari inequality. Let us define the Bers embedding by
$\beta([\mu])=\mathcal{S}(w^{\mu}|_{\mathbb{D}^{*}})\in A_{\infty}(\mathbb{D}^{*}).$
The Bers embedding is a holomorphic map of complex Banach manifolds.
The Banach space of bounded harmonic Beltrami differential on $\mathbb{D}$ is defined by
$\Omega^{-1,1}(\mathbb{D})=\{\,\mu\in L^{\infty}(\mathbb{D})\,|\,\mu=e^{-\psi}\bar{\phi},\,\phi\in A_{\infty}(\mathbb{D})\,\}$
where $e^{\psi}$ denotes the hyperbolic density function on $\mathbb{D}$ and $A_{\infty}(\mathbb{D})$ is defined as $A_{\infty}(\mathbb{D}^{*})$. The decomposition
(2.7) $L^{\infty}(\mathbb{D})=\Omega^{-1,1}(\mathbb{D})\oplus\mathcal{N}(\mathbb{D})$
identifies the holomorphic tangent space $T_{[0]}T(1)=L^{\infty}(\mathbb{D})/\mathcal{N}(\mathbb{D})$ at the origin in $T(1)$ as follows
(2.8) $T_{[0]}T(1)\cong\Omega^{-1,1}(\mathbb{D}).$
The complex linear mapping $D_{0}\beta$ induces the isomorphism $\Omega^{-1,1}(\mathbb{D})\cong A_{\infty}(\mathbb{D}^{*})$ of the holomorphic tangent spaces to $T(1)$ and $A_{\infty}(\mathbb{D}^{*})$ at the origin.
The unit ball $L^{\infty}(\mathbb{D})_{1}$ carries a group structure induced by the composition of quasiconformal maps. The group law $\lambda=\nu\star\mu^{-1}$ is defined through $w_{\lambda}=w_{\nu}\circ w_{\mu}^{-1}$. The explicit formula of the group law is given by
(2.9) $w_{\mu}^{*}(\lambda):=\lambda\circ w_{\mu}\frac{\overline{\partial_{z}w_{\mu}}}{{\partial_{z}w_{\mu}}}=\frac{\nu-\mu}{1-\nu\bar{\mu}}\,.$
For $\mu\in L^{\infty}(\mathbb{D})_{1}$, using the group structure of $L^{\infty}(\mathbb{D})_{1}$, we define the right translation $R_{\mu}$ on $L^{\infty}(\mathbb{D})_{1}$. The induced right translations on $T(1)$
(2.10) $R_{[\mu]}:T(1)\longrightarrow T(1),\qquad[\lambda]\mapsto[\lambda\star\mu]$
are biholomorphic automorphisms of $T(1)$. Therefore, $D_{0}R_{[\mu]}:T_{[0]}T(1)\to T_{[\mu]}T(1)$ is a complex linear isomorphism between holomorphic tangent spaces, and we obtain the identification of $T_{[\mu]}T(1)\cong\Omega^{-1,1}(\mathbb{D})$.
For $\mu\in L^{\infty}(\mathbb{D})_{1}$, let $U_{\mu}$ be the image of the ball of radius 2 in $A_{\infty}(\mathbb{D}^{*})$ under the map $h_{\mu}^{-1}=\mathrm{P}\circ R_{\mu}\circ\Lambda$ where $\Lambda$ is the inverse of $D_{0}\beta$. Then the maps
$h_{\mu\nu}:=h_{\mu}\circ h_{\nu}^{-1}:h_{\nu}(U_{\mu}\cap U_{\nu})\to h_{\mu}(U_{\mu}\cap U_{\nu})$
are biholomorphic as functions on the Banach space $A_{\infty}(\mathbb{D}^{*})$. The structure of $T(1)$ as a complex Banach manifold modeled on the Banach space $A_{\infty}(\mathbb{D}^{*})$ is explicitly described by the complex-analytic atlas given by the open covering
$T(1)=\bigcup_{\mu\in L^{\infty}(\mathbb{D})_{1}}U_{\mu}$
with coordinate maps $h_{\mu}$ and the transition maps $h_{\mu\nu}$. Complex coordinates on $T(1)$ defined by the coordinate charts $(U_{\mu},h_{\mu})$ are called Bers coordinates. For every $\nu\in\Omega^{-1,1}(\mathbb{D})$, set $\phi=D_{0}\beta(\nu)$ and define a holomorphic vector field $\frac{\partial}{\partial\epsilon_{\nu}}$ on $U_{0}$ by setting
$Dh_{0}\Big{(}\frac{\partial}{\partial\epsilon_{\nu}}\Big{)}=\phi$
at all points in $U_{0}$. At every point $[\mu]\in U_{0}$, identified with the corresponding harmonic Beltrami differential $\mu$, the vector field $\frac{\partial}{\partial\epsilon_{\nu}}$ in terms of the Bers coordinates of $U_{\mu}$ correspond to
$\tilde{\phi}=D_{\mu}h_{\mu}\Big{(}\frac{\partial}{\partial\epsilon_{\nu}}\Big{)}=\Big{(}D_{\mu}h_{\mu}(D_{\mu}h_{0})^{-1}\Big{)}(\phi)=D_{0}(\beta\circ\mathrm{P})(D_{\mu}R_{\mu}^{-1}(\Lambda(\phi))).$
Using identification $\Omega^{-1,1}(\mathbb{D})\cong A_{\infty}(\mathbb{D}^{*})$, provided by $D_{0}\beta$,
(2.11) $\frac{\partial}{\partial\epsilon_{\nu}}\Big{|}_{\mu}=D_{0}\mathrm{P}(D_{\mu}R_{\mu}^{-1}(\nu))=D_{0}\mathrm{P}(R(\nu,\mu)),$
where
(2.12) $R(\nu,\mu):=D_{\mu}R_{\mu}^{-1}(\nu)=\Big{(}\frac{\nu}{1-|\mu|^{2}}\frac{(w_{\mu})_{z}}{(\bar{w}_{\mu})_{\bar{z}}}\Big{)}\circ w_{\mu}^{-1}.$
2.2. Weil-Petersson universal Teichmüller space
Let us consider
$A_{2}(\mathbb{D}^{*})=\Big{\{}\,\phi:\mathbb{D}^{*}\to\mathbb{C}\,|\,\text{holomorphic},\int_{\mathbb{D}^{*}}|\phi|^{2}e^{-\psi}\,d^{2}z<\infty\,\Big{\}}\subset A_{\infty}(\mathbb{D}^{*})$
where $d^{2}z=dxdy$ for $z=x+\sqrt{-1}y$. Denote by $\mathcal{O}(\mathbb{D})_{1}$ the subgroup of $L^{\infty}(\mathbb{D})_{1}$ generated by $\mu\in\Omega^{-1,1}(\mathbb{D})$ with $||\mu||_{\infty}<\delta$ where $\delta$ is a postive real number satisfying condition in Corollary 2.6 in [15]. For every $\mu\in\mathcal{O}(\mathbb{D})_{1}$, let $V_{\mu}\subset U_{\mu}\subset T(1)$ be the image under the map $h_{\mu}^{-1}=\mathrm{P}\circ R_{\mu}\circ\Lambda$ of the open ball of radius $\sqrt{\pi/3}$ about the origin in $A_{2}(\mathbb{D}^{*})$. Let $\tilde{h}_{\mu}=h_{\mu}\Big{|}_{V_{\mu}}:V_{\mu}\to A_{2}(\mathbb{D^{*}})$. Now we consider the covering
$T(1)=\bigcup_{\mu\in\mathcal{O}(\mathbb{D})_{1}}V_{\mu}$
with the coordinate maps $\tilde{h}_{\mu}:V_{\mu}\to A_{2}(\mathbb{D}^{*})$ and the transition maps
$\tilde{h}_{\mu\nu}=\tilde{h}_{\mu}\circ\tilde{h}_{\nu}^{-1}:\tilde{h}_{\nu}(V_{\mu}\cap V_{\nu})\to\tilde{h}_{\mu}(V_{\mu}\cap V_{\nu}).$
By Theorem 2.10 in [15], the above covering endows $T(1)$ with the structure of a complex Hilbert manifold modeled on the Hilbert space $A_{2}(\mathbb{D}^{*})$. But, the space $T(1)$ is not connected with respect to the topology induced by the Hilbert manifold structure.
For the Hilbert space of harmonic Beltrami differentials on $\mathbb{D}$,
$H^{-1,1}(\mathbb{D})=\{\,\mu\in L^{\infty}(\mathbb{D})\,|\,\mu=e^{-\psi}\bar{\phi},\,\phi\in A_{2}(\mathbb{D})\,\}\subset\Omega^{-1,1}(\mathbb{D}),$
and for $[\mu]\in T(1)$, let $D_{0}R_{[\mu]}(H^{-1,1}(\mathbb{D}))$ be the subspace of the tangent space $T_{[\mu]}T(1)=D_{0}R_{[\mu]}(\Omega^{-1,1}(\mathbb{D}))$, which is equipped with a Hilbert space structure isomorphic to $H^{-1,1}(\mathbb{D})$. Let $\mathcal{D}_{T}$ be the distribution on $T(1)$, defined by the assignment
$T(1)\ni[\mu]\mapsto D_{0}R_{[\mu]}(H^{-1,1}(\mathbb{D}))\subset T_{[\mu]}T(1).$
By Theorem 2.3 in [15], for every $[\mu]\in T(1)$, the linear mapping
$D_{0}(\beta\circ R_{[\mu]}):H^{-1,1}(\mathbb{D})\to A_{2}(\mathbb{D}^{*})$
is a topological isomorphism. Moreover, by Theorem 2.13 in [15], the Bers embedding
$\beta:T(1)\to\beta(T(1))\subset A_{\infty}(\mathbb{D}^{*})$
is a biholomorphic mapping of Hilbert manifolds. Hence the distribution $\mathcal{D}_{T}$ on $T(1)$ is integrable. Their integral manifolds correspond to the components $(\phi+A_{2}(\mathbb{D}^{*}))\cap\beta(T(1))$. For every $[\mu]\in T(1)$, we denote by $T_{[\mu]}(1)$ the component of the Hilbert manifold $T(1)$ containing $[\mu]$. The Hilbert manifold $T_{[\mu]}(1)$ is the integral manifold of the distribution $\mathcal{D}_{T}$ passing through $[\mu]\in T(1)$. In particular, the component of the origin $0\in T(1)$ is denoted by $T_{0}(1)$, and is called Weil-Petersson universal Teichmüller space.
The Weil-Petersson metric on the Hilbert manifold $T_{0}(1)$ is a Hermitian metric defined by the Hilbert space inner product on tangent space, which is identified with the Hilbert space $H^{-1,1}(\mathbb{D})$ by a right translation. Thus the Weil-Petersson metric is a right invariant metric on $T_{0}(1)$ defined at the origin of $T_{0}(1)$ by
(2.13) $\langle\mu,\nu\rangle=\int_{\mathbb{D}}\mu\bar{\nu}\,e^{\psi}\,d^{2}z,\qquad\text{for}\quad\mu,\nu\in H^{-1,1}(\mathbb{D})=T_{0}T_{0}(1).$
To every $\mu\in H^{-1,1}(\mathbb{D})$, there corresponds a vector field $\frac{\partial}{\partial\epsilon_{\mu}}$ over $V_{0}$. We set for every $\kappa\in V_{0}$,
$g_{\mu\bar{\nu}}(\kappa)=\Big{\langle}\frac{\partial}{\partial\epsilon_{\mu}}\Big{|}_{\kappa},\frac{\partial}{\partial\epsilon_{\nu}}\Big{|}_{\kappa}\Big{\rangle}_{WP}=\int_{\mathbb{D}}D_{0}\mathrm{P}(R(\mu,\kappa))\overline{D_{0}\mathrm{P}(R(\nu,\kappa))}\,e^{\psi}\,d^{2}z$
where $R(\mu,\kappa)$ is given in (2.12). The Weil-Petersson metric extends to other charts $V_{\mu}$ by right translations.
3. Maximal discs of Weil-Petersson class
In this section, we introduce and study maximal discs of Weil-Petersson class in the 3-dimensional Anti de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. The deformation space of maximal discs of Weil-Petersson class is understood as the cotangent bundle $T^{*}T_{0}(1)$ of the Weil-Petersson universal Teichmüller space $T_{0}(1)$, and we show that this can be identified with $T_{0}(1)\times T_{0}(1)$ by the map originally defined by Mess in [10]. We refer to the appendix of this paper for the basic terminology related to $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ and to [5] for the extensive introduction for the Anti de Sitter geometry.
3.1. Maximal discs and Gauss maps
For an Beltrami differential $\mu$ representing an element in the universal Teichmüller space $T(1)$, let us denote by $(\mathbb{D},z)$ the corresponding unit disc with the coordinate map $z$. Note that $z$ is given by the quasi-conformal map $w_{\mu}:\mathbb{D}\to\mathbb{D}$ introduced in the subsection 2.1. Let us also fix the conformal structure on the standard disc $\mathbb{D}\subset\mathbb{C}$ and denote it by $(\mathbb{D},w)$ with its coordinate map $w$.
For a given $(\mathbb{D},z)$, let us denote the spacelike conformal embedding of $(\mathbb{D},z)$ into the 3-dimensional anti de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ by
$\sigma:(\mathbb{D},z)\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}.$
For a spacelike conformal embedding $\sigma$, the image $\Sigma:=\sigma(\mathbb{D})$ is called maximal if the mean curvature $H_{\sigma}$ identically vanishes on $\Sigma$.
By Proposition 3.1 and Theorem 3.3 of [2], For the maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, each component of the Gauss map $G_{\pm}:\Sigma\to(\mathbb{D},w)$ is a harmonic map, that is, satisfies the following equality
(3.1) $(G_{\pm})_{z\bar{z}}+(\psi_{w}\circ G_{\pm})(G_{\pm})_{z}(G_{\pm})_{\bar{z}}=0,$
where $e^{\psi(w)}|dw|^{2}$ denotes the hyperbolic metric on $(\mathbb{D},w)$. The first fundamental form $I$, that is, the induced metric on $\Sigma$ is given by
(3.2) $I=e^{\phi}|dz|^{2}:=e^{\psi\circ G_{\pm}}\big{|}(G_{\pm})_{z}\big{|}^{2}|dz|^{2}.$
The Hopf differential $\Phi(G_{\pm})$ of the harmonic map $G_{\pm}:\Sigma\to(\mathbb{D},w)$ satisfies
(3.3) $\Phi(G_{+})=e^{\psi\circ G_{+}}(G_{+})_{z}(\overline{G_{+}})_{z}\,dz^{2}=-\Phi(G_{-})=-e^{\psi\circ G_{-}}(G_{-})_{z}(\overline{G_{-}})_{z}\,dz^{2},$
and the second fundamental forms of $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is given by
(3.4) $II=\frac{\sqrt{-1}}{2}\big{(}\Phi(G_{+})-\overline{\Phi(G_{+})}\big{)}.$
We refer to the section 3 of [2] and the section 5 of [7] for more details about these equalities.
By the Gauss equation for $\Sigma$, we have
(3.5) $2\phi_{z\bar{z}}=e^{\phi}-e^{-\phi}|\Phi(G_{\pm})|^{2}.$
Note that (3.5) can also be derived from the harmonic map equation for $G_{\pm}$ given in (3.1). We also have
Proposition 3.1.
The Gaussian curvature $K_{\phi}$ for the metric $e^{\phi}|dz|^{2}$ on $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is given by
(3.6) $K_{\phi}:=-2\phi_{z\bar{z}}e^{-\phi}=-1+|\mu_{G_{\pm}}|^{2}.$
Proof.
From (3.2) and (3.1), denoting by $G=G_{\pm}$,
$\phi_{z}=(\psi_{w}\circ G)G_{z}+\frac{G_{zz}}{G_{z}},$
and
$\phi_{z\bar{z}}=(\psi_{w\bar{w}}\circ G)\big{(}|G_{z}|^{2}-|G_{\bar{z}}|^{2}\big{)}=\frac{1}{2}e^{\psi\circ G}\big{(}|G_{z}|^{2}-|G_{\bar{z}}|^{2}\big{)}.$
Here the second equality follows from the Liouville equation for $e^{\psi}$, that is, $\psi_{w\bar{w}}=1/2e^{\psi}$. Hence, we have
$K_{\phi}=-2\phi_{z\bar{z}}e^{-\phi}=-e^{\psi\circ G}\big{(}|G_{z}|^{2}-|G_{\bar{z}}|^{2}\big{)}\cdot e^{-\psi\circ G}|G_{z}|^{-2}=-1+|\mu_{G}|^{2}.$
By the Gauss equation for $\Sigma$, $K_{\phi}=-1+\kappa^{2}$ where ±$\pm\kappa$ denote the principal curvatures of the maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. Hence,
(3.7) $|\mu_{G_{\pm}}|^{2}=\kappa^{2}.$
For the hyperbolic metric on $(\mathbb{D},w)$, its pullback by $G_{\pm}$ determines a new conformal structure on $\Sigma$ respectively. We denote them by $(\mathbb{D},z_{\pm})$ respectively. By Theorem 3.2 of [18], for the Hopf differential ±$\pm\Phi=\Phi(G_{\pm})$ on $\Sigma\cong(\mathbb{D},z)$ and $(\mathbb{D},z_{\pm})$, there is a harmonic map $F_{\pm}:(\mathbb{D},z)\to(\mathbb{D},z_{\pm})$ such that the associated Hopf differential is ±$\pm\Phi$. Moreover, we have
Proposition 3.2.
(3.8) $e^{\phi}=e^{\psi\circ G_{\pm}}|(G_{\pm})_{z}|^{2}=e^{\psi\circ F_{\pm}}|(F_{\pm})_{z}|^{2},\qquad\mu_{G_{\pm}}=\mu_{F_{\pm}},$
where $e^{\psi}$ denotes the hyperbolic metric density on $(\mathbb{D},w)$, $(\mathbb{D},z_{\pm})$ respectively.
Proof.
Recall that the harmonic map condition of $G_{\pm}$, $F_{\pm}$ depends on the conformal structure on $\Sigma$ rather than the metic on $\Sigma$, and that the concerned quantities of $G_{\pm}$, $F_{\pm}$ are the square of the holomorphic energy densities times the hyperbolic metric density on $\Sigma$, that is,
$e^{\psi\circ G_{\pm}}|(G_{\pm})_{z}|^{2}=|\partial G_{\pm}|^{2}e^{\psi(z)},\qquad e^{\psi\circ F_{\pm}}|(F_{\pm})_{z}|^{2}=|\partial F_{\pm}|^{2}e^{\psi(z)}.$
Then, by Theorem 2.1 of [18] (see also the section 4 of [22]), the functions
(3.9) $\mathfrak{w}(G_{\pm}):=1/2\log|\partial G_{\pm}|^{2},\qquad\mathfrak{w}(F_{\pm}):=1/2\log|\partial F_{\pm}|^{2}$
satisfy the following equation
$14e^{-\psi}\mathfrak{w}_{z\bar{z}}=-(K\circ u)\big{(}e^{2\mathfrak{w}}-e^{-2\mathfrak{w}}e^{-2\psi}|\Phi|^{2}\big{)}-$ 1
where $K$ denotes the Gaussian curvature of the target disc of $G_{\pm}$ or $F_{\pm}$ and the map $u$ is $G_{\pm}$ or $F_{\pm}$ depending on the case. But, $K(u)\equiv-1$ for these cases. Then, by Corollary 8 of [22], the solution of this equation with $K\circ u\equiv-1$ is unique. Hence, $\mathfrak{w}(G_{+})=\mathfrak{w}(G_{-})=\mathfrak{w}(F_{+})=\mathfrak{w}(F_{-})$, which implies the first equality of $e^{\psi\circ G_{\pm}}|(G_{\pm})_{z}|^{2}=e^{\psi\circ F_{\pm}}|(F_{\pm})_{z}|^{2}$ in (3.8). Note that the Beltrami differentials $\mu_{G_{\pm}}$ and $\mu_{F_{\pm}}$ satisfy
$\mu_{G_{\pm}}=\pm\overline{\Phi}e^{-\phi}=\mu_{F_{\pm}},$
whhich implies the second equality in (3.8). ∎
By Proposition 3.2, $(F_{+},F_{-})$ shares some basic properties with the Gauss map $(G_{+},G_{-})$, and we call $(F_{+},F_{-})$ the induced Gauss map.
Combining the above constructions for two cases of $G_{\pm}$, we have the following commutative diagram
[…]
Here $z,z_{+},z_{-}$ denote the quasi-conformal maps from $(\mathbb{D},w)$ to $(\mathbb{D},z)$, $(\mathbb{D},z_{+})$, $(\mathbb{D},z_{-})$ respectively. We also abuse notations by using the same notations $F_{+}$, $F_{-}$ for the central horizontal maps, which are obviously induced by $F_{\pm}:\Sigma\to(\mathbb{D},z_{\pm})$. We may also assume that all the maps $z,z_{\pm}$ and $F_{\pm}$ are normalized, that is, preserves $1,-1,-\sqrt{-1}$.
Remark 3.3.
By the construction of the harmonic map $F_{\pm}:(\mathbb{D},z)\to(\mathbb{D},z_{\pm})$, we have $F_{\pm}\circ z=z_{\pm}$ as explained above. Hence, the harmonic map $F_{\pm}$ is compatible with the coordinate maps $z,z_{\pm}$ in the above sense. But, such a property does not holds for $G_{\pm}:(\mathbb{D},z)\to(\mathbb{D},w)$, that is, $G_{\pm}\circ z\neq\mathrm{id}$.
By Theorem 4.1 of [2], Theorem 4.1 of [4], and the previous construction, we have
Proposition 3.4.
There are one-to-one correspondences between the followings:
  1. (1)
    a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ such that the Gaussian curvature of the induced metric is negative,
  2. (2)
    an orientation preserving minimal diffeomorphism $G:(\mathbb{D},w)\to(\mathbb{D},w)$ where $(\mathbb{D},w)$ is equipped with the hyperbolic metric,
  3. (3)
    an orientation preserving minimal diffeomorphism $F:(\mathbb{D},z_{+})\to(\mathbb{D},z_{-})$ where $(\mathbb{D},z_{\pm})$ is equipped with the hyperbolic metric,
  4. (4)
    a quasisymmetric homeomorphism $h:\mathbb{S}^{1}\to\mathbb{S}^{1}$.
Some remarks are in order about Proposition 3.4. For $(1)\to(2)$, the orientation preserving minimal diffeomorphism $G:\mathbb{D}\to\mathbb{D}$ is obtained by putting $G=G_{+}\circ G_{-}^{-1}$ for the Gauss map $(G_{+},G_{-}):\Sigma\to\mathbb{D}\times\mathbb{D}$. For $(2)\to(4)$, the quasisymmetric homeomorphism $h:\mathbb{S}^{1}\to\mathbb{S}^{1}$ is obtained by restricting $G$ to the boundary $\partial{\mathbb{D}}$. For $(4)\to(1)$, the maximal disc $\Sigma$ is obtained such that $\partial{\Sigma}$ is the graph of $h:\mathbb{S}^{1}\to\mathbb{S}^{1}$. For the equivalence of $(1),(3),(4)$, one can use a similar construction by (3.10) as above instead of $(2)$.
Remark 3.5.
By Proposition 3.4, the followings equivalently represent the same points in the universal Teichmüller space $T(1)$ respectively.
  1. (1)
    the conformal structure of $\mathbb{D}_{\pm}:=(\mathbb{D},z_{\pm})$,
  2. (2)
    the quasisymmetric homeomorphism $z_{\pm}|_{\mathbb{S}^{1}}=(F_{\pm}\circ z)|_{\mathbb{S}^{1}}$,
  3. (3)
    the Beltrami differential $\mu_{z_{\pm}}$ of the map $z_{\pm}$.
For a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, the induced conformal structure on $\Sigma$ determines $(\mathbb{D},z)$, hence the Beltrami differential $\mu$ of $z$. The Hopf differentials of the induced Gauss maps $F_{\pm}$ are ±$\pm\Phi$ by the construction. Hence, the maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ determines a pair $(\mu,\Phi)$ in $T^{*}T(1)$. Now, we introduce the following definition,
Definition 3.6.
For a maximal conformal embedding $\sigma$ of $(\mathbb{D},z)$, the maximal disc ${\Sigma}\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is of Weil-Petersson class if the associate pair $(\mu,\Phi)\in T^{*}T_{0}(1)$.
Proposition 3.7.
For a maximal disc of Weil-Petersson class $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ with the induced metric $e^{\phi}|dz|^{2}$ and the induced Gauss map $(F_{+},F_{-}):\Sigma\to\mathbb{D}_{+}\times\mathbb{D}_{-}$, we have
  1. (1)
    the total curvature is finite, that is,
    $\int_{\Sigma}\kappa^{2}e^{\phi}\ d^{2}z=\int_{\Sigma}|\mu_{G_{\pm}}|^{2}e^{\phi}\,d^{2}z=\int_{\Sigma}|\mu_{F_{\pm}}|^{2}e^{\phi}\,d^{2}z<\infty,$
  2. (2)
    the anti-holomorphic energies of $G_{\pm}$ and $F_{\pm}$ are finite respectively, that is,
    $\int_{\Sigma}|\Phi(G_{\pm})|^{2}e^{-\phi}\,d^{2}z=\int_{\Sigma}|\Phi(F_{\pm})|^{2}e^{-\phi}\,d^{2}z<\infty.$
Proof.
By the condition that $\Sigma$ is Weil-Petersson class and Lemma 3.3 of [15],
$\int_{\Sigma}|\mu_{F_{\pm}}|^{2}e^{\psi}\,d^{2}z<\infty.$
This implies that $|\mu_{F_{\pm}}|(z)\to 0$ as $|z|\to 1$. Now, recall that there exists a nonzero constant $a$ such that the curvature $-1<K_{\phi}<-a^{2}$ by (3.6) and $|\mu_{\pm}|\to 0$. Then, we have that $e^{\phi}<a^{-2}e^{\psi}$ by Lemma 4.9 of [21]. By these facts and $\kappa^{2}=|\mu_{F_{\pm}}|^{2}$,
(3.11) $\int_{\Sigma}\kappa^{2}e^{\phi}\ d^{2}z=\int_{\Sigma}|\mu_{F_{\pm}}|^{2}e^{\phi}\,d^{2}z<a^{-2}\int_{\Sigma}|\mu_{F_{\pm}}|^{2}e^{\psi}\,d^{2}z<\infty$
This completes the proof of the item (1). Then the item (2) follows from this easily noticing that $|\Phi(F_{\pm})|^{2}e^{-\phi}=|\mu_{F_{\pm}}|^{2}e^{\phi}$.
3.2. Mess map
For a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, its first and second fundamental forms have the local expressions in terms of the induced conformal structure on $\Sigma$ and the Hopf differential $\Phi(G_{\pm})$ as in (3.2) and (3.4). Then, each component of the Gauss map $G_{\pm}$ determines the conformal structures $(\mathbb{D},z_{\pm})$ respectively. Denoting by $I_{\pm}$ the hyperbolic metric on $(\mathbb{D},z_{\pm})$, we have
(3.12) $I_{\pm}=I(E\pm JI^{-1}II\cdot\,,E\pm JI^{-1}II\cdot\,)$
where $E$ denotes the identity map and $J$ denotes the complex structure on $T\Sigma$. The equality (3.12) is proved in Lemma 3.16 of [8]. Note that the equality (3.12) makes sense originally for the pair $(I,II)$ given by $(\mu,\Phi)$ as in (3.2) and (3.4), which is associated to a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. In particular, the Beltrami differential $\mu$ corresponds to the induced conformal structure on $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$.
Naturally, we can extend the above correspondence given in (3.12) for any $(\mu,\Phi)\in T^{*}T(1)$ where $\mu$ may not correspond to the induced conformal structure on $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. For a $(\mu,\Phi)\in T^{*}T(1)$, we solve the Gauss equation (3.5) with respect to $\Phi$ on $(\mathbb{D},z)$ where $z$ denotes the coordinate corresponding to $\mu$. Then we define by $I(\mu,\Phi):=e^{\phi}|dz|^{2}$ where $\phi$ denotes the solution of the Gauss equation. Moreover, by Theorem 3.2 of [18], there exist harmonic maps
$F_{\pm}:(\mathbb{D},z)\to(\mathbb{D},z_{\pm}),$
which we also call induced Gauss map in this general case. Now, we introduce the Mess map defined by
(3.13) $\mathrm{Mess}:T^{*}T(1)\to T(1)\times T(1),$
which maps $(\mu,\Phi)$ to $(I_{+},I_{-})$ using (3.12) for $I(\mu,\Phi)$ and $II$ given by (3.4). By the result in the section 5 of [7], the map $\mathrm{Mess}$ is bijective.
Remark 3.8.
By the above constructions of the induced Gauss map $F_{\pm}$, one can see that the one-to-one correspondence of the items (3) and (4) in Proposition 3.4 still holds in this general situation.
Remark 3.9.
For a totally geodesic maximal disc $(\mathbb{D},w)\cong\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, we can see that $I_{\pm}=I=e^{\psi(w)}|dw|^{2}$ by (3.12) since $II=0$. On the other hand, for a pair $(\mu,0)\in T^{*}T(1)$ representing a different conformal structure $(\mathbb{D},z)$ on a (real) totally geodesic maximal disc $\Sigma$, the map $\mathrm{Mess}$ sends $(\mu,0)\in T^{*}T(1)$ to $I_{\pm}=e^{\psi(z)}|dz|^{2}$, which is not the induced metric $e^{\psi(w)}|dw|^{2}$. Hence, the image of the induced Gauss map $(F_{+},F_{-})$ of the totally geodesic disc $(\mathbb{D},z)$ is always realized as the diagonal minimal Lagrangian submanifold in $(\mathbb{D},z_{+})\times(\mathbb{D},z_{-})$ where both discs are equipped with the hyperbolic metrics. This also implies that the map $\mathrm{Mess}$ sends the image of zero section of $T^{*}T(1)$ to the diagonal set in $T(1)\times T(1)$.
Restricting the Mess map to the holomorphic cotangent bundle of the Weil-Petersson Teichmüller space $T_{0}(1)$, we have
Theorem 3.10.
The following restriction of $\mathrm{Mess}$ is a diffeomorphism,
(3.14) $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1).$
Proof.
Recall that a point in $T^{*}T_{0}(1)$ is represented by a pair $(\mu,\Phi)$ for a Beltrami differential $\mu$ on $\mathbb{D}$ and a holomorphic quadratic differential $\Phi$ on $\mathbb{D}$ satisfying the conditions
(3.15) $\int_{\mathbb{D}}|\mu|^{2}e^{\psi}\,d^{2}w<\infty,\qquad\int_{\mathbb{D}}|\Phi|^{2}e^{-\psi}\,d^{2}z<\infty.$
Here $(\mathbb{D},w)$ denotes the origin in $T_{0}(1)$ and $(\mathbb{D},z)$ represents the conformal structure on $\Sigma$ determined by $\mu$. Note that the first inequality follows from Lemma 3.3 of [15] and the second inequality follows from the definition of $T^{*}T_{0}(1)$. Hence, these imply that the restrictions of maps $F_{+},F_{-},z$ to the boundary are quasisymmetric homeomorphisms of $\mathbb{S}^{1}$ representing points in $T_{0}(1)$ respectively. Since the composition of two quasisymmetric homeomorphisms representing elements in $T_{0}(1)$ is a quasisymmetric homeomorphism in $T_{0}(1)$, the restriction of $z_{\pm}=F_{\pm}\circ z$ to the boundary $\mathbb{D}$ is a quasisymmetric homeomorphism in $T_{0}(1)$. Hence, one can conclude that $\mathrm{Mess}$ maps $T^{*}T_{0}(1)$ into $T_{0}(1)\times T_{0}(1)$.
For given quasisymmetric homeomorphisms $f_{\pm}:\mathbb{S}^{1}\to\mathbb{S}^{1}$ representing two points in $T_{0}(1)$, let us take quasi-conformal extension determined by the corresponding the Beltrami differentials and we denote them by $z_{\pm}:(\mathbb{D},w)\to(\mathbb{D},z_{\pm})$. Now, by Proposition 3.4 and Remark 3.8, for the pair $(f_{+},f_{-})$, there exists a minimal Lagrangian diffeomorphism extension $F:(\mathbb{D},z_{-})\to(\mathbb{D},z_{+})$ of $f_{+}\circ f_{-}^{-1}:\mathbb{S}^{1}\to\mathbb{S}^{1}$. Moreover, for $(\mathbb{D},z)$ denoting the conformal structure of the graph of $F=F_{+}\circ F_{-}^{-1}$ induced from $(\mathbb{D},z_{+})\times(\mathbb{D},z_{-})$, there exist harmonic maps $F_{\pm}:(\mathbb{D},z)\to(\mathbb{D},z_{\pm})$ such that $F|_{\partial{\mathbb{D}}}=f_{+}\circ f_{-}^{-1}$. Here note that $(\mathbb{D},z)$ may not be same as the conformal structure of $\Sigma$ induced from $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. By the above constructions and Lemma 3.3 in [15], we have
(3.16) $\int_{\mathbb{D}}|\mu_{F}|^{2}e^{\psi}\,d^{2}z_{-}<\infty.$
By (2.9), the Beltrami differential $\mu_{F}$ has the following expression
(3.17) $\mu_{F}\circ F_{-}\frac{\overline{(F_{-})_{z}}}{(F_{-})_{z}}=\frac{\mu_{+}-\mu_{-}}{1-\mu_{+}\overline{\mu}_{-}}=\frac{2\mu_{+}}{1+|\mu_{+}|^{2}},$
where $\mu_{\pm}=\mu_{F_{\pm}}$ so that
(3.18) $\left|\mu_{F}\circ F_{-}\right|=\frac{2|\mu_{+}|}{1+|\mu_{+}|^{2}}=\frac{2|\mu_{-}|}{1+|\mu_{-}|^{2}}.$
Combining these,
(3.19) $\int_{\mathbb{D}}|\mu_{F}|^{2}e^{\psi}\,d^{2}z_{-}=\,\int_{\Sigma}\left(|\mu_{F}|^{2}\circ F_{-}\right)\,e^{\psi\circ F_{-}}\,\left(\left|(F_{-})_{z}\right|^{2}-\left|(F_{-})_{\bar{z}}\right|^{2}\right)\,d^{2}z\\ =\,\int_{\Sigma}\left(|\mu_{F}|^{2}\circ F_{-}\right)\,e^{\psi\circ F_{-}}\,\left|(F_{-})_{z}\right|^{2}\left(1-\left|\mu_{-}\right|^{2}\right)\,d^{2}z\\ =\,\int_{\Sigma}\frac{4|\mu_{-}|^{2}}{(1+|\mu_{-}|^{2})^{2}}e^{\phi}\left(1-\left|\mu_{-}\right|^{2}\right)\,d^{2}z.$
Since $[\mu_{F}]\in T_{0}(1)$, $|\mu_{F}(z)|\to 0$ as $|z|\to 1$, so that $|\mu_{-}(z)|\to 0$ as $z$ approaches to $\partial\Sigma$ by (3.18). Hence, there exists $C>0$ such that
(3.20) $C<\frac{1-|\mu_{-}|^{2}}{(1+|\mu_{-}|^{2})^{2}}<1\qquad\text{over}\quad\Sigma.$
Hence, by (3.19) and (3.20),
(3.21) $\int_{\mathbb{D}}|\mu_{F}|^{2}e^{\psi}\,d^{2}z<\infty\qquad\text{ if and only if}\qquad\int_{\Sigma}{|\mu_{\pm}|^{2}}e^{\psi}\,d^{2}z\leq\int_{\Sigma}{|\mu_{\pm}|^{2}}e^{\phi}\,d^{2}z<\infty.$
This means that $F_{\pm}|_{\partial{\mathbb{D}}}$ represents points in $T_{0}(1)$. For the given map $f_{\pm}$, which is the same as $z_{\pm}|_{\partial{\mathbb{D}}}$, representing points in $T_{0}(1)$ respectively, we have that $z|_{\partial{\mathbb{D}}}=(F_{\pm}^{-1}\circ z_{\pm})|_{\partial{\mathbb{D}}}$ represent a point in $T_{0}(1)$. Hence, the pair $(\mu_{z},\Phi)$ represents a point in $T^{*}T_{0}(1)$ where $\mu_{z}$ is the Beltrami differential of $z$ and ±$\pm\Phi$ denotes the Hopf differential of $F_{\pm}$. Here the fact $\Phi\in T^{*}T_{0}(1)$ follows from (3.21). In conclusion, the map $\mathrm{Mess}$ maps $(\mu_{z},\Phi)\in T^{*}T_{0}(1)$ to the given $(z_{+},z_{-})$. Hence, the map $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is onto.
To prove the injectivity of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$, we consider the following equalities
(3.22) $\mu_{z_{\pm}}=\frac{\mu_{z}\pm z^{*}(\mu_{+})}{1\pm\bar{\mu}_{z}\,z^{*}(\mu_{+})}$
where $\mu_{+}=\mu_{F_{+}}$. If the map $\mathrm{Mess}$ is not injective, there are two pairs $(a,b)$ and $(a^{\prime},b^{\prime})$ for $(\mu_{z},z^{*}(\mu_{F}))$ satisfying the above equalities for given $\mu_{z_{\pm}}$. Assuming this, we can derive the following equalities
(3.23) $a\pm b+\bar{a^{\prime}}bb^{\prime}\pm a\bar{a^{\prime}}b^{\prime}=a^{\prime}\pm b^{\prime}+\bar{a}bb^{\prime}\pm\bar{a}a^{\prime}b.$
Combing these, we have
(3.24) $a-a^{\prime}=(\bar{a}-\bar{a^{\prime}})bb^{\prime}.$
Hence, if $a\neq a^{\prime}$, we have $|bb^{\prime}|=1$. But, this contradicts to $|bb^{\prime}|<1$. Hence, we must have $a=a^{\prime}$. This also implies that $b=b^{\prime}$ by (3.23). This completes the proof of the injectivity of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$.
To complete the proof of the claim, it is enough to show that the differential of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is an isomorphism by the inverse function theorem. This will be proved in Proposition 4.8. ∎
As in the proof of Theorem 3.10, one can also prove
Theorem 3.11.
For any $(\mu,\Phi)\in T^{*}T_{0}(1)$, the anti-holomorphic energy of the associated induced harmonic maps $F_{\pm}$ defines a finite valued functional over $T^{*}T_{0}(1)$.
4. Symplectic structure on $T^{*}T_{0}(1)$
In this section, we first derive some variational formulae for the deformations of a maximal disc in the 3-dimensional Anti de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. Using these formulae, we prove the relationship of symplectic forms in terms of the map $\mathrm{Mess}$.
4.1. Variational formulae
For a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, we denote by $\Sigma^{\epsilon}\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ its deformation family for a small real parameter $\epsilon$. In general, such a deformation consists of two parts: one is a deformation of conformal structures on the maximal disc and the deformation of the way of embedding of the maximal disc in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. Here we allow the conformal structure of the maximal disc, which needs not to be the same as the induced conformal structure determined by the embedding. For a deformation of conformal structure of the maximal disc, we denote it by $\Sigma^{\epsilon}\cong(\mathbb{D},z_{\epsilon})$. The other part of a deformation of $\Sigma^{\epsilon}$, which is determined by the way of embedding, can be described in terms of a deformation of the induced Gauss map $(F_{+},F_{-}):\Sigma\to(\mathbb{D},z_{+})\times(\mathbb{D},z_{-})$.
To deal with such a general situation of a deformation of a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, we consider the following diagram
[…]
Here $f^{\epsilon}$and $h^{\epsilon}$ denote quasiconformal maps with the corresponding harmonic Beltrami differential $\epsilon\nu_{f}$ and $\epsilon\nu_{h}$ respectively. Hence, the quasi-conformal maps $f^{\epsilon}$, $h^{\epsilon}$ satisfy the following Beltrami differential equations
(4.2) $f^{\epsilon}_{\bar{z}}=\epsilon\nu_{f}f^{\epsilon}_{z},\qquad h^{\epsilon}_{\bar{u}}=\epsilon\nu_{h}h^{\epsilon}_{u}.$
Taking derivative at $\epsilon=0$, we have
(4.3) $\dot{f}_{\bar{z}}=\nu_{f},\qquad\dot{h}_{\bar{u}}=\nu_{h}.$
Lemma 4.1.
For $\nu_{h}=\dot{h}_{\bar{z}}$ satisfying $h^{\epsilon}\circ F=H^{\epsilon}$,
(4.4) $\nu_{h}=R(\dot{\mu}_{H},\mu_{F})=\Big{(}\frac{\dot{\mu}_{H}}{1-|\mu_{F}|^{2}}\frac{F_{z}}{\overline{F}_{\bar{z}}}\Big{)}\circ{F}^{-1}.$
Here $\dot{\mu}_{H}=\frac{d}{d\epsilon}\big{|}_{\epsilon=0}\mu_{H^{\epsilon}}$.
Proof.
From (2.9), we derive the following expression for the Beltrami differential $\mu_{H^{\epsilon}}$ of $H^{\epsilon}$
(4.5) $\mu_{H^{\epsilon}}=\frac{\mu_{F}+F^{*}(\epsilon\nu_{h})}{1+\bar{\mu}_{F}F^{*}(\epsilon\nu_{h})}.$
Taking derivative of the equality (4.5) at $\epsilon=0$, we have $\dot{\mu}_{H}=(1-|\mu_{F}|^{2})F^{*}(\nu_{h})$. From this, the equality (4.4) follows easily. ∎
For a family of quasiconformal maps $f^{\epsilon}$ with $f^{\epsilon}_{\bar{z}}=\epsilon\nu f^{\epsilon}_{z}$ and a smooth family of tensors $\omega^{\epsilon}$ of type $(\ell,m)$, set
$(f^{\epsilon})^{*}(w^{\epsilon})=\omega^{\epsilon}\circ f^{\epsilon}\big{(}(f^{\epsilon})_{z}\big{)}^{\ell}\big{(}(\overline{f^{\epsilon}})_{\bar{z}}\big{)}^{m}.$
The Lie derivatives of the family $\omega^{\epsilon}$ along a vector field $\frac{\partial}{\partial\epsilon_{\nu}}$ is defined by
$L_{\nu}\omega=\frac{\partial}{\partial\epsilon}\Big{|}_{\epsilon=0}(f^{\epsilon})^{*}(w^{\epsilon}).$
Proposition 4.2.
If the diagram (4.1) is commutative, that is, $F^{\epsilon}\circ f^{\epsilon}=H^{\epsilon}=h^{\epsilon}\circ F$,
(4.6) $L_{\nu_{f}}\mu_{F}=(1-|\mu_{F}|^{2})F^{*}(\nu_{h})-(\nu_{f}-\bar{\nu}_{f}\mu_{F}^{2}).$
Proof.
From the commutative diagram in (4.1), we have
(4.7) $\frac{\mu_{F}+\epsilon F^{*}(\nu_{h})}{1+\epsilon\bar{\mu}_{F}F^{*}(\nu_{h})}=\frac{\epsilon\nu_{f}+(f^{\epsilon})^{*}(\mu_{F}^{\epsilon})}{1+\epsilon\bar{\nu}_{f}(f^{\epsilon})^{*}(\mu_{F}^{\epsilon})}.$
Taking derivative $\epsilon=0$ for (4.7), we have
(4.8) $(1-|\mu_{F}|^{2})F^{*}(\nu_{h})=(\nu_{f}+L_{\nu_{f}}\mu_{F})-\bar{\nu}_{f}\mu_{F}^{2}.$
This completes the proof. ∎
Proposition 4.3.
If the diagram (4.1) is commutative and $\nu_{f}$ is a harmonic Beltrami differential, then
(4.9) $\dot{\mu}_{H}=\nu_{f}-\bar{\nu}_{f}\mu_{F}^{2}+L_{\nu_{f}}\mu_{F}.$
Proof.
The equality (4.9) easily follows from (4.4) and (4.6). ∎
Proposition 4.4.
For a family of Hopf differentials $\Phi^{\epsilon}=e^{\psi^{\epsilon}\circ F^{\epsilon}}F^{\epsilon}_{z}\overline{F}^{\epsilon}_{z}$ on $\Sigma^{\epsilon}$ satisfying the commutative diagram (4.1),
(4.10) $\begin{split}L_{{\nu}_{f}}\,\Phi=&\,\Phi\big{(}F^{*}({{\nu_{h}}})\bar{\mu}_{F}-{\bar{\nu}_{f}}\mu_{F}\big{)}+\overline{\Phi}\big{(}F^{*}(\bar{{\nu}}_{h})\mu_{F}^{-1}-\bar{{\nu}}_{f}\mu_{F}^{-1}\big{)}\\ =&\,e^{\phi}\big{(}F^{*}({{\nu_{h}}})\bar{\mu}_{F}^{2}-{\bar{\nu}_{f}}|\mu_{F}|^{2}\big{)}+e^{\phi}\big{(}F^{*}(\bar{{\nu}}_{h})-\bar{{\nu}}_{f}\big{)}.\end{split}$
Proof.
From the definition of Lie derivative, we have
(4.11) $L_{{\nu}_{f}}\,\Phi=\frac{\partial}{\partial\epsilon}\Big{|}_{\epsilon=0}\Big{(}e^{\psi^{\epsilon}\circ F^{\epsilon}\circ f^{\epsilon}}F^{\epsilon}_{z}\circ f^{\epsilon}\,\overline{F}_{{z}}^{\epsilon}\circ f^{\epsilon}\,f^{\epsilon}_{z}{f}^{\epsilon}_{{z}}\Big{)}.$
From $h^{\epsilon}\circ F=F^{\epsilon}\circ f^{\epsilon}$, we have
(4.12) $(\dot{h}\circ F)_{z}=\frac{\partial}{\partial\epsilon}\Big{|}_{\epsilon=0}\Big{(}F^{\epsilon}_{z}\circ f^{\epsilon}f^{\epsilon}_{z}\Big{)}+F_{\bar{z}}\dot{\bar{f}}_{z}.$
Combining (4.11) and (4.12),
$L_{{\nu_{f}}}\,\Phi=$ $\ e^{\psi\circ F}\Big{(}\dot{\psi}+\psi_{u}\dot{h}+\psi_{\bar{u}}\dot{\bar{h}}\Big{)}\circ FF_{z}\overline{F}_{{z}}$
$+e^{\psi\circ F}\Big{(}(\dot{h}\circ F)_{z}-F_{\bar{z}}\dot{\bar{f}}_{z}\Big{)}\overline{F}_{{z}}+e^{\psi\circ F}F_{z}\Big{(}(\dot{\bar{h}}\circ F)_{\bar{z}}-\overline{F}_{\bar{z}}\dot{\bar{f}}_{{z}}\Big{)}$
$=$ $\ e^{\psi\circ F}\Big{(}\dot{\psi}+\psi_{u}\dot{h}+\psi_{\bar{u}}\dot{\bar{h}}+\dot{h}_{z}+\dot{\bar{h}}_{\bar{z}}\Big{)}\circ FF_{z}\overline{F}_{{z}}$
$+e^{\psi\circ F}\Big{(}\dot{h}_{\bar{z}}\circ F\overline{F}_{z}\overline{F}_{{z}}-\dot{\bar{f}}_{z}F_{\bar{z}}\overline{F}_{{z}}+\dot{\bar{h}}_{{z}}\circ F{F}_{{z}}{F}_{{z}}-\dot{\bar{f}}_{{z}}F_{{z}}\overline{F}_{\bar{z}}\Big{)}$
$=$ $\ e^{\psi\circ F}F_{z}\overline{F}_{z}\Big{(}\dot{h}_{\bar{z}}\circ F\overline{F}_{\bar{z}}\overline{F}_{z}(F_{z}\overline{F}_{\bar{z}})^{-1}-\dot{\bar{f}}_{{z}}F_{\bar{z}}F_{z}^{-1}\Big{)}$
$+e^{\psi\circ F}F_{\bar{z}}\overline{F}_{\bar{z}}\Big{(}\dot{\bar{h}}_{z}\circ FF_{z}^{2}(F_{\bar{z}}\overline{F}_{\bar{z}})^{-1}-\dot{\bar{f}}_{z}F_{z}F_{\bar{z}}^{-1}\Big{)}$
$=$ $\ \Phi\big{(}F^{*}({{\nu_{h}}})\bar{\mu}_{F}-{\bar{\nu}_{f}}\mu_{F}\big{)}+\overline{\Phi}\big{(}F^{*}(\bar{{\nu}}_{h})\mu_{F}^{-1}-\bar{{\nu}}_{f}\mu_{F}^{-1}\big{)}.$
For the third equality above, we used the following equality for a variation of the hyperbolic metrics
(4.13) $\dot{\psi}+\psi_{u}\dot{h}+\psi_{\bar{u}}\dot{\bar{h}}+\dot{h}_{z}+\dot{\bar{h}}_{\bar{z}}=0,$
which follows from the Ahlfors’ lemma given in [1]. This completes the proof. ∎
Remark 4.5.
Expressions of the Lie derivatives in (4.6) and (4.10) are given by both the harmonic Beltrami differentials on $\Sigma$ and the pullback by $F$ of the harmonic Beltrami differentials on the image of $F$. This is due to the deformation of (4.1) involve the deformations on the source and target discs of the harmonic map $F$ simultaneously.
For a family of harmonic diffeomorphism $F_{\pm}^{\epsilon}:\Sigma^{\epsilon}\to\mathbb{D}^{\epsilon}_{\pm}:=(\mathbb{D},(z_{\pm})_{\epsilon})$, let us consider the following commutative diagram
[…]
such that $h^{\epsilon}_{\pm}\circ F_{\pm}=F^{\epsilon}_{\pm}\circ f^{\epsilon}$. Moreover, we also assume the corresponding equalities to (4.2) and (4.3) for $f^{\epsilon}$ and $h^{\epsilon}_{\pm}$.
Proposition 4.6.
For a family of Hopf differentials $\Phi^{\epsilon}=\Phi^{\epsilon}_{+}$ of $F^{\epsilon}_{+}$ satisfying the commutative diagram (4.14),
(4.15) $L_{\nu_{f}}\Phi=\frac{1}{2}e^{\phi}\Big{(}(F^{*}_{+}(\nu_{+})-F^{*}_{-}(\nu_{-}))\bar{\mu}_{F}^{2}+(F^{*}_{+}(\bar{\nu}_{+})-F^{*}_{-}(\bar{\nu}_{-}))\Big{)},$
(4.16) $\nu_{f}=\frac{1}{2}(1+|\mu_{F}|^{2})^{-1}\Big{(}(F^{*}_{+}(\nu_{+})+F^{*}_{-}(\nu_{-}))+(F^{*}_{+}(\bar{\nu}_{+})+F^{*}_{-}(\bar{\nu}_{-}))\mu_{F}^{2}\Big{)},$
where $\nu_{\pm}$ denotes the Beltrami differential $\nu_{h_{\pm}}$ and $\mu_{F}=\mu_{F_{+}}$.
Proof.
By $\Phi_{\pm}=e^{\phi}\bar{\mu}_{F_{\pm}}$ and Proposition 4.4 applied to $F^{\epsilon}_{\pm}$,
(4.17) $L_{{\nu}_{f}}\,\Phi_{\pm}=e^{\phi}\big{(}F_{\pm}^{*}({{\nu_{\pm}}})\bar{\mu}_{F_{\pm}}^{2}-{\bar{\nu}_{f}}|\mu_{F_{\pm}}|^{2}\big{)}+e^{\phi}\big{(}F^{*}_{\pm}(\bar{{\nu}}_{\pm})-\bar{{\nu}}_{f}\big{)}.$
The difference of the equalities (4.17) for $\Phi_{\pm}$ with $\Phi_{+}=-\Phi_{-}$ gives the equality (4.15).
To prove (4.16), we apply Proposition 4.2 to $F^{\epsilon}_{\pm}$ and take the sum of them. Then we have
(4.18) $2\nu_{f}-2\bar{\nu}_{f}\mu_{F}^{2}=(1-|\mu_{F}|^{2})\big{(}F^{*}_{+}(\nu_{+})+F^{*}_{-}(\nu_{-})\big{)}.$
Combining this and its conjugated equality and noting $\overline{F^{*}_{\pm}(\nu_{\pm})}=F^{*}_{\pm}(\bar{\nu}_{\pm})$ gives the equality (4.16). ∎
Remark 4.7.
In the equalities (4.15) and (4.16), the pullback of the Beltrami differentials $\nu_{\pm}$, $\bar{\nu}_{\pm}$ by $F_{\pm}$ and $\mu_{F_{\pm}}$ have the same tensor type $(-1,1)$. But their dependence are different and they play the different roles in the following sense. The Beltrami differential $\mu_{F_{\pm}}$ depends only on $F_{\pm}:\Sigma\to(\mathbb{D},z_{\pm})$ and play the role of a function over $T_{0}(1)$. On the other hand, the pullbacks of $\nu_{\pm}$, $\bar{\nu}_{\pm}$ by $F_{\pm}$ depend on the deformation and play the role of vectors in the tangent bundle of $T_{0}(1)$.
4.2. Symplectic forms
For a harmonic diffeomorphism $F_{\pm}:\Sigma\to\mathbb{D}_{\pm}:=(\mathbb{D},z_{\pm})$ in the diagram (4.14), we denote by $(\mu_{+},\mu_{-})$ the Beltrami differential representing $(\mathbb{D},z_{\pm})$ respectively. As explained in the subsection 2.2, let us fix an orthonormal basis of in $H^{-1,1}(\mathbb{D})$ and we obtain the corresponding orthonormal basis for $T_{[\mu]}T_{0}(1)$ by the right translations, that is, given by $D_{0}R_{[\mu]}(H^{-1,1}(\mathbb{D}))$ for any $[\mu]\in T_{0}(1)$. We denote by $\{\nu_{\pm,k}\}_{k\in\mathbb{Z}}$ the orthonormal basis of the tangent space $T_{[\mu_{\pm}]}T_{0}(1)$ obtained in this way respectively. For a harmonic Beltrami differential $\nu_{\pm,k}$ on $(\mathbb{D},z_{\pm})$, we solve the Beltrami differential equation
(4.19) $(h^{\epsilon}_{\pm})_{\bar{z}}=\epsilon\nu_{\pm,k}(h^{\epsilon}_{\pm})_{z}\qquad\text{where}\quad z=z_{\pm}$
to obtain a quasi-conformal map $h^{\epsilon}_{\pm}:\mathbb{D}_{\pm}\to\mathbb{D}_{\pm}^{\epsilon}$. Note that $h^{\epsilon}_{\pm}$ and $\mathbb{D}^{\epsilon}_{\pm}$ depends on $k$, but we do not denote it for simplicity of notations. Now, for given two above discs $\mathbb{D}^{\epsilon}_{\pm}$, as the way of the inverse of the map $\mathrm{Mess}$, we construct a maximal disc $\Sigma^{\epsilon}$ in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ such that there exist harmonic maps $F_{\pm}^{\epsilon}:\Sigma^{\epsilon}\to\mathbb{D}^{\epsilon}_{\pm}$. Then, there exists a unique quasi-conformal map $f^{\epsilon}:\Sigma\to\Sigma^{\epsilon}$ such that
(4.20) $h^{\epsilon}_{+}\circ F_{+}=F^{\epsilon}_{+}\circ f^{\epsilon},\qquad h^{\epsilon}_{-}\circ F_{-}=F^{\epsilon}_{-}\circ f^{\epsilon}.$
This can be easily follows from the fact that the maximal disc $\Sigma^{\epsilon}$ is the graph of the map $F^{\epsilon}_{+}\circ(F^{\epsilon}_{-})^{-1}$, and that a point in $\Sigma^{\epsilon}$ is determined by its two coordinates given by $F^{\epsilon}_{\pm}$. Note that the map $f^{\epsilon}$ and the maximal disc $\Sigma^{\epsilon}$ depend on $k$. Now we put $\nu_{k}$ to be the Beltrami differential on $\Sigma$ such that $(f^{\epsilon}_{k})_{\bar{z}}=\epsilon\nu_{k}(f^{\epsilon}_{k})_{z}$. Here $f^{\epsilon}_{k}$ denotes the quasi-conformal map $f^{\epsilon}$ corresponding $k$.
Now we define the variation of $\mu$ along $\nu_{k}$ by
(4.21) $\delta_{k}\mu:={\nu_{k}}=\frac{1}{2}(1+|\mu_{F}|^{2})^{-1}\Big{(}(F^{*}_{+}(\nu_{+,k})+F^{*}_{-}(\nu_{-,k}))+(F^{*}_{+}(\bar{\nu}_{+,k})+F^{*}_{-}(\bar{\nu}_{-,k}))\mu_{F}^{2}\Big{)}$
where the second equality follows from (4.16). By (4.6), we can see that $F^{*}_{\pm}(\nu_{\pm,k})$ is determined by $\nu_{k}$ and $\mu_{F}=\mu_{F_{+}}$. We also define the variation of $\Phi$ along $\nu_{k}$ by
(4.22) $\delta_{k}\Phi:=L_{\nu_{k}}\Phi=\frac{1}{2}e^{\phi}\Big{(}(F^{*}_{+}(\nu_{+,k})-F^{*}_{-}(\nu_{-,k}))\bar{\mu}_{F}^{2}+(F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}(\bar{\nu}_{-,k}))\Big{)}$
where the second equality follows from (4.15).
Proposition 4.8.
The differential of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is an isomorphism.
Proof.
By the proof of Theorem 3.10, the map $\mathrm{Mess}$ is one-to-one and onto, hence its inverse $\mathrm{Mess}^{-1}$ is well-defined. Now, we want to show that the differential of $\mathrm{Mess}^{-1}$ is an isomorphism. First, we observe that the differential of $\mathrm{Mess}^{-1}$ is given by (4.21) and (4.22). In order to analyze these maps, we decompose a tangent space $T(T_{0}(1)\times T_{0}(1))$ at $z_{\pm}|_{\mathbb{S}^{1}}$ into the diagonal tangent spaces and the anti-diagonal tangent space. Then, along the diagonal tangent space, we have $F^{*}_{+}(\nu_{+,k})=F^{*}_{-}(\nu_{-,k})$. Hence, the component of the differential of $\mathrm{Mess}^{-1}$ given in (4.22) vanishes over this subspace. If the component of the differential of $\mathrm{Mess}^{-1}$ given in (4.21) has a nontrivial kernel, then we should have
$F^{*}_{+}(\nu_{+})+F^{*}_{+}(\bar{\nu}_{+})\mu_{F}^{2}=0$
for a harmonic Beltrami differential $\nu_{+}$. Combining this with its conjugated equality, we have $F^{*}_{+}(\nu_{+})(1-|\mu_{F}|^{4})=0$, which can not happen since $|\mu_{F}|<1$ and $F_{+}$ is a diffeomorphism. Hence, the component of the differential of $\mathrm{Mess}^{-1}$ given in (4.21) is injective. To prove that this component is onto over the diagonal tangent space, pick a harmonic Beltrami differential $\nu$ representing a tangent vector $\delta\mu$, then there exists a Beltrami differential $\nu_{+}$ such that
$\nu-\bar{\nu}\mu_{F}^{2}=(1-|\mu_{F}|^{2})F_{+}^{*}(\nu_{+})$
by the fact $F$ is a diffeomorphism. This implies that
$\nu=(1+|\mu_{F}|^{2})^{-1}\Big{(}F^{*}_{+}(\nu_{+})+F^{*}_{+}(\bar{\nu}_{+})\mu_{F}^{2}\Big{)}.$
Since $\nu$ can be written in terms of $\nu_{k}$’s, the same is true for $\nu_{+}$ in terms of $\nu_{+,k}$’s. Hence, the component is onto over the diagonal tangent space. In a similar way, we can show that the differential of $\mathrm{Mess}^{-1}$ given in (4.22) is bijective over anti-diagonal tangent space. Hence, these together means that the differential of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is an isomorphism. This completes the proof. ∎
By Proposition 4.8, we can see that $\{\delta_{k}\mu\}_{k\in\mathbb{Z}}$ and $\{\delta_{k}\Phi\}_{k\in\mathbb{Z}}$ together are basis of the tangent bundle of $T^{*}T_{0}(1)$ at $(\mu,\Phi)$. Using this fact, we define the symplectic form $\omega_{\mathrm{C}}$ on the cotangent bundle $T^{*}T_{0}(1)$ over an open neighborhood $V_{0}$ of the origin in $T_{0}(1)$ by
(4.23) $\omega_{\mathrm{C}}=\int_{\Sigma}\Big{(}\sum_{k\in\mathbb{Z}}\sqrt{-1}\delta_{k}\Phi\wedge\delta_{k}\mu-\sqrt{-1}\delta_{k}\overline{\Phi}\wedge\delta_{k}\bar{\mu}\Big{)}\,d^{2}z.$
Note that the wedge product $\wedge$ is taken for the vectors $F^{*}_{\pm}(\nu_{\pm,k})$ which are vector terms on the right hand side of (4.21) and (4.22). This formula defines the symplectic form on the coordinate chart $V_{0}$. This symplectic form also extends to other charts $V_{\mu}$ by right translation. The definition of $w_{\mathrm{C}}$ involves the choice of basis $\{\nu_{\pm,k}|\,k\in\mathbb{Z}\}$ a priori. But, we will show that $\omega_{\mathrm{C}}$ does not depend on this choice (see Remark 4.11).
Proposition 4.9.
The following equality holds
(4.24) $\begin{split}\omega_{\mathrm{C}}=&-\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F_{+}}|^{2})\,e^{\phi}\\ &\qquad\qquad\cdot\,\sum_{k\in\mathbb{Z}}\Big{(}F^{*}_{+}(\nu_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}(\nu_{-,k})\wedge F^{*}_{-}(\bar{\nu}_{-,k})\Big{)}\,d^{2}z.\end{split}$
Proof.
By (4.23), (4.15), and (4.16), recalling $\mu_{F}=\mu_{F_{+}}$,
$\sqrt{-1}\Big{(}\sum_{k\in\mathbb{Z}}\delta_{k}\Phi\wedge\delta_{k}\mu-\delta_{k}\overline{\Phi}\wedge\delta_{k}\bar{\mu}\Big{)}$
$=$ $\frac{\sqrt{-1}}{4}(1+|\mu_{F}|^{2})^{-1}\,e^{\phi}\,\sum_{k\in\mathbb{Z}}\Big{(}(F^{*}_{+}(\nu_{+,k})-F^{*}_{-}(\nu_{-,k}))\bar{\mu}_{F}^{2}+(F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}(\bar{\nu}_{-,k})\Big{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\wedge\Big{(}(F^{*}_{+}(\nu_{+,k})+F^{*}_{-}(\nu_{-,k}))+(F^{*}_{+}(\bar{\nu}_{+,k})+F^{*}_{-}(\bar{\nu}_{-,k})\mu_{F}^{2}\Big{)}$
$-\frac{\sqrt{-1}}{4}(1+|\mu_{F}|^{2})^{-1}\,e^{\phi}\,\sum_{k\in\mathbb{Z}}\Big{(}(F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}(\bar{\nu}_{-,k})){\mu}_{F}^{2}+(F^{*}_{+}({\nu}_{+,k})-F^{*}_{-}({\nu}_{-,k})\Big{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\quad\wedge\Big{(}(F^{*}_{+}(\bar{\nu}_{+,k})+F^{*}_{-}(\bar{\nu}_{-,k}))+(F^{*}_{+}({\nu}_{+,k})+F^{*}_{-}({\nu}_{-,k})\bar{\mu}_{F}^{2}\Big{)}$
$=$ $\frac{\sqrt{-1}}{4}(1+|\mu_{F}|^{2})^{-1}\,e^{\phi}\,\Big{[}\sum_{k\in\mathbb{Z}}(F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}(\bar{\nu}_{-,k})\wedge(F^{*}_{+}({\nu}_{+,k})+F^{*}_{-}({\nu}_{-,k}))$
$\qquad\qquad\qquad\qquad\qquad+(F^{*}_{+}(\nu_{+,k})-F^{*}_{-}(\nu_{-,k}))\wedge(F^{*}_{+}(\bar{\nu}_{+,k})+F^{*}_{-}(\bar{\nu}_{-,k})|\mu_{F}|^{4}$
$\qquad\qquad\qquad\qquad\qquad-(F^{*}_{+}({\nu}_{+,k})-F^{*}_{-}({\nu}_{-,k}))\wedge(F^{*}_{+}(\bar{\nu}_{+,k})+F^{*}_{-}(\bar{\nu}_{-,k}))$
$\qquad\qquad\qquad\qquad\qquad-(F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}(\bar{\nu}_{-,k}))\wedge(F^{*}_{+}({\nu}_{+,k})+F^{*}_{-}({\nu}_{-,k})|{\mu}_{F}|^{4}\Big{]}$
$=$ $\frac{\sqrt{-1}}{4}(1-|\mu_{F}|^{2})\,e^{\phi}\,\Big{[}\sum_{k\in\mathbb{Z}}(F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}(\bar{\nu}_{-,k})\wedge(F^{*}_{+}({\nu}_{+,k})+F^{*}_{-}({\nu}_{-,k}))$
$\qquad\qquad\qquad\qquad\qquad-(F^{*}_{+}({\nu}_{+,k})-F^{*}_{-}({\nu}_{-,k}))\wedge(F^{*}_{+}(\bar{\nu}_{+,k})+F^{*}_{-}(\bar{\nu}_{-,k}))\Big{]}$
$=$ $-\frac{\sqrt{-1}}{2}(1-|\mu_{F}|^{2})\,e^{\phi}\,\Big{[}\sum_{k\in\mathbb{Z}}F^{*}_{+}({\nu}_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}({\nu}_{-,k})\wedge F^{*}_{-}(\bar{\nu}_{-,k})\Big{]}.$
This completes the proof. ∎
Recall that the basis $\{\nu_{\pm,k}\}_{k\in\mathbb{Z}}$ is an orthonormal basis at $T^{*}_{[\mu_{\pm}]}T_{0}(1)$ and that the Weil-Petersson symplectic form $\omega_{\mathrm{WP}}$ on $T^{*}_{[\mu_{\pm}]}T_{0}(1)$ has the following expression,
(4.25) $\omega_{\mathrm{WP}}=\frac{\sqrt{-1}}{2}\int_{\mathbb{D}}e^{\psi}\big{(}\sum_{k\in\mathbb{Z}}\nu_{\pm,k}\wedge\bar{\nu}_{\pm,k}\big{)}\,d^{2}z.$
Theorem 4.10.
The map $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is a symplectic diffeomorphism such that
(4.26) $\omega_{\mathrm{C}}=-\mathrm{Mess}^{*}_{+}(\omega_{\mathrm{WP}})+\mathrm{Mess}^{*}_{-}(\omega_{\mathrm{WP}}).$
Here $\mathrm{Mess}_{\pm}:=\pi_{\pm}\circ\mathrm{Mess}$ where $\pi_{\pm}$ denotes the projection map from $T_{0}(1)\times T_{0}(1)$ onto the first or second factor respectively.
Proof.
By Proposition 4.9,
(4.27) $\begin{split}\omega_{\mathrm{C}}=&-\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F}|^{2})\,e^{\phi}\\ &\qquad\qquad\quad\cdot\Big{[}\sum_{k\in\mathbb{Z}}F^{*}_{+}({\nu}_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}({\nu}_{-,k})\wedge F^{*}_{-}(\bar{\nu}_{-,k})\Big{]}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\int_{\Sigma}\,e^{\psi\circ F_{\pm}}\big{(}|(F_{\pm})_{z}|^{2}-|(F_{\pm})_{\bar{z}}|^{2}\big{)}\\ &\qquad\qquad\quad\cdot\Big{[}\sum_{k\in\mathbb{Z}}F^{*}_{+}({\nu}_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})-F^{*}_{-}({\nu}_{-,k})\wedge F^{*}_{-}(\bar{\nu}_{-,k})\Big{]}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\Big{[}\mathrm{Mess}^{*}_{+}\int_{\mathbb{D}}e^{\psi}\,\Big{(}\sum_{k\in\mathbb{Z}}\nu_{+,k}\wedge\bar{\nu}_{+,k}\Big{)}\,d^{2}z\\ &\qquad\qquad\qquad\qquad\qquad-\mathrm{Mess}^{*}_{-}\int_{\mathbb{D}}e^{\psi}\Big{(}\sum_{k\in\mathbb{Z}}\nu_{-,k}\wedge\bar{\nu}_{-,k}\Big{)}\,d^{2}z\Big{]}.\end{split}$
Hence, the above equalities imply
$\omega_{\mathrm{C}}=-\mathrm{Mess}^{*}_{+}(\omega_{\mathrm{WP}})+\mathrm{Mess}^{*}_{-}(\omega_{\mathrm{WP}}).$
This completes the proof. ∎
Remark 4.11.
By Theorem 4.10, we can see that the symplectic form $\omega_{\mathrm{C}}$ does not depend on the choice of the basis $\{\nu_{\pm,k}:\,k\in\mathbb{Z}\}$.
Remark 4.12.
By the proof of Theorem 4.10, we can see that the pullback $\mathrm{Mess}_{\pm}^{*}(\omega_{\mathrm{WP}})$ has the following expression
(4.28) $\mathrm{Mess}^{*}_{\pm}(\omega_{\mathrm{WP}})=\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F}|^{2})\,e^{\phi}\,\Big{(}\sum_{k\in\mathbb{Z}}F^{*}_{\pm}(\nu_{\pm,k})\wedge F^{*}_{\pm}(\bar{\nu}_{\pm,k})\Big{)}\,d^{2}z.$
The closedness of this pullback 2-forms on $T^{*}T_{0}(1)$ just follows from the commutativity of $d$-differential and pullback map. But, it can be also proved as in the arguments in Lemma 2.7 of [24] or Theorem 7.4 of [15], whose original argument comes from [1]. For this argument, we need the fact that the Lie derivative of the density $(1-|\mu_{F}|^{2})e^{\phi}$ vanishes, which follows from Propositions 5.1 and 5.2.
Corollary 4.13.
For the symplectic diffeomorphism $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$,
  1. (1)
    $\mathrm{Mess}$ maps the Lagrangian subspace given by the zero section of $T^{*}T_{0}(1)$ to the diagonal subset in $T_{0}(1)\times T_{0}(1)$,
  2. (2)
    $\mathrm{Mess}$ maps the Lagrangian subspace given by the fiber $T^{*}_{0}T_{0}(1)$ to the anti-diagonal subset in $T_{0}(1)\times T_{0}(1)$.
Proof.
From the bottom part of the diagram (3.10), we have that $z_{\pm}=F_{\pm}\circ z$ in general. For the case of the Lagrangian subspace of the zero section of $T^{*}T_{0}(1)$, the map $F_{\pm}$ becomes trivial since the Hopf differential $\Phi$ is trivial. Hence, the disc $(\mathbb{D},z_{\pm})$ coincides with $(\mathbb{D},z)$, that is, $z_{\pm}=z$. In conclusion, $\mathrm{Mess}$ maps the Lagrangian of the zero section to the diagonal subset in $T_{0}(1)\times T_{0}(1)$. For the case of the Lagrangian subspace of the fiber $T_{0}^{*}T_{0}(1)$, the disc $(\mathbb{D},z)$ coincides with $(\mathbb{D},w)$, that is, $z_{\pm}=F_{\pm}$. Hence, the Beltrami differential of $z_{\pm}$ is the same as the Beltrami differential $\mu_{F_{\pm}}$. Moreover, $\mu_{F_{\pm}}=\pm\overline{\Phi}e^{-\phi}$. In conclusion, $\mathrm{Mess}$ maps the Lagrangian space of fiber $T_{0}^{*}T_{0}(1)$ to the anti-diagonal subset in $T_{0}(1)\times T_{0}(1)$. ∎
5. Variation of anti-holomorphic energy
In this section, we prove that the anti-holomorphic energy is a potential function of the symplectic form $\omega_{C}$ restricted to the subspaces $T_{0}(1)^{\pm}$ of $T^{*}T_{0}(1)$. Recall that $T_{0}(1)^{+}$, $T_{0}(1)^{-}$ denote the inverse image of $\mathrm{Mess}$ of $T_{0}(1)\times\{0\}$, $\{0\}\times T_{0}(1)$ respectively.
First, we consider a variation for the holomorphic energy density and anti-holomorphic energy density.
Proposition 5.1.
For a family of anti-holomorphic energy densities $e^{\psi^{\epsilon}\circ F^{\epsilon}}|F^{\epsilon}_{\bar{z}}|^{2}$ on $\Sigma^{\epsilon}$ satisfying the commutative diagram (4.1),
(5.1) $L_{{\nu}_{f}}\big{(}\,e^{\psi\circ F}|F_{\bar{z}}|^{2}\,\big{)}=\Phi\big{(}F^{*}({{\nu_{h}}})-{\nu_{f}}\big{)}+\overline{\Phi}\big{(}F^{*}(\bar{{\nu}}_{h})-\bar{{\nu}}_{f}\big{)}.$
Proof.
From the definition of Lie derivative, we have
(5.2) $L_{{\nu}}\,\big{(}\,e^{\psi\circ F}|F_{\bar{z}}|^{2}\,\big{)}=\frac{\partial}{\partial\epsilon}\Big{|}_{\epsilon=0}\Big{(}e^{\psi^{\epsilon}\circ F^{\epsilon}\circ f^{\epsilon}}F^{\epsilon}_{\bar{z}}\circ f^{\epsilon}\,\overline{F}_{{z}}^{\epsilon}\circ f^{\epsilon}\,f^{\epsilon}_{z}\bar{f}^{\epsilon}_{\bar{z}}\Big{)}.$
From $h^{\epsilon}\circ F=F^{\epsilon}\circ f^{\epsilon}$, we have
(5.3) $\frac{\partial}{\partial\epsilon}\Big{|}_{\epsilon=0}\Big{(}F^{\epsilon}_{\bar{z}}\circ f^{\epsilon}\bar{f}^{\epsilon}_{\bar{z}}\Big{)}=(\dot{h}\circ F)_{\bar{z}}-F_{{z}}\dot{{f}}_{\bar{z}},\qquad\frac{\partial}{\partial\epsilon}\Big{|}_{\epsilon=0}\Big{(}\overline{F}^{\epsilon}_{{z}}\circ f^{\epsilon}{f}^{\epsilon}_{{z}}\Big{)}=(\dot{\bar{h}}\circ F)_{{z}}-\overline{F}_{\bar{z}}\dot{\bar{f}}_{{z}}$
Combining (5.2) and (5.3),
$L_{{\nu_{f}}}\,\big{(}\,e^{\psi\circ F}|F_{\bar{z}}|^{2}\,\big{)}=$ $\ e^{\psi\circ F}\Big{(}\dot{\psi}+\psi_{u}\dot{h}+\psi_{\bar{u}}\dot{\bar{h}}\Big{)}\circ FF_{\bar{z}}\overline{F}_{{z}}$
$+e^{\psi\circ F}\Big{(}(\dot{h}\circ F)_{\bar{z}}-F_{{z}}\dot{{f}}_{\bar{z}}\Big{)}\overline{F}_{{z}}+e^{\psi\circ F}F_{\bar{z}}\Big{(}(\dot{\bar{h}}\circ F)_{{z}}-\overline{F}_{\bar{z}}\dot{\bar{f}}_{{z}}\Big{)}$
$=$ $\ e^{\psi\circ F}\Big{(}\dot{\psi}+\psi_{u}\dot{h}+\psi_{\bar{u}}\dot{\bar{h}}+\dot{h}_{z}+\dot{\bar{h}}_{\bar{z}}\Big{)}\circ FF_{\bar{z}}\overline{F}_{{z}}$
$+e^{\psi\circ F}\Big{(}\dot{h}_{\bar{z}}\circ F\overline{F}_{\bar{z}}\overline{F}_{{z}}-\dot{{f}}_{\bar{z}}F_{{z}}\overline{F}_{{z}}+\dot{\bar{h}}_{{z}}\circ F{F}_{{z}}{F}_{\bar{z}}-\dot{\bar{f}}_{{z}}F_{\bar{z}}\overline{F}_{\bar{z}}\Big{)}$
$=$ $\ e^{\psi\circ F}F_{z}\overline{F}_{z}\Big{(}\dot{h}_{\bar{z}}\circ F\overline{F}_{\bar{z}}F_{z}^{-1}-\dot{f}_{\bar{z}}\Big{)}+e^{\psi\circ F}F_{\bar{z}}\overline{F}_{\bar{z}}\Big{(}\dot{\bar{h}}_{z}\circ FF_{z}\overline{F}_{\bar{z}}^{-1}-\dot{\bar{f}}_{z}\Big{)}$
$=$ $\ \Phi\Big{(}F^{*}({\nu}_{h})-{\nu}_{f}\Big{)}+\overline{\Phi}\Big{(}F^{*}({\bar{\nu}}_{h})-{\bar{\nu}_{f}}\Big{)}.$
For the third equality above, we used the equality (4.13) for a variation of the hyperbolic metrics as before. This completes the proof. ∎
In the same way as the above, we can prove
Proposition 5.2.
For a family of holomorphic energy densities $e^{\phi^{\epsilon}}=e^{\psi^{\epsilon}\circ F^{\epsilon}}|F^{\epsilon}_{z}|^{2}$ on $\Sigma^{\epsilon}$ satisfying the commutative diagram (4.1),
(5.4) $L_{{\nu}_{f}}\,e^{\phi}=\Phi\big{(}F^{*}({{\nu_{h}}})-{\nu_{f}}\big{)}+\overline{\Phi}\big{(}F^{*}(\bar{{\nu}}_{h})-\bar{{\nu}}_{f}\big{)}.$
Recall that $T_{0}(1)^{\pm}$ is a symplectic space with the restriction of $\omega_{\mathrm{C}}$, which is the same as $\mp\mathrm{Mess}^{*}(\omega_{\mathrm{WP}})$ by Theorem 4.10. Using (2.9), along $T_{0}(1)^{\pm}$, we can derive
(5.5) $\mu_{z_{\pm}}=\frac{2\mu_{z}}{1+|\mu_{z}|^{2}},\qquad\mu_{F_{+}}=-(z^{-1})^{*}(\mu_{z})$
where $\mu_{z_{\pm}}$, $\mu_{z}$ denotes the Beltrami differential of the map $z_{\pm}$, $z$ from $(\mathbb{D},w)$.
Proposition 5.3.
Under the condition $F_{+}^{*}(\nu_{+})=0$ or $F_{-}^{*}(\nu_{-})=0$,
(5.6) $\frac{1}{2}\big{(}L_{\nu}-iL_{i\nu}\big{)}\big{(}e^{\psi\circ F}|F_{\bar{z}}|^{2}\big{)}=e^{\phi}\bar{\mu}_{F}\nu.$
Proof.
We will prove only for the case $F_{-}^{*}(\nu_{-})=0$ since the proof for the other case is the same. By (5.1),
(5.7) $L_{{\nu}}\big{(}e^{\psi\circ F}|F_{\bar{z}}|^{2}\big{)}=\Phi\big{(}F^{*}({{\nu_{h}}})-{\nu}\big{)}+\overline{\Phi}\big{(}F^{*}({\bar{\nu}_{h}})-\bar{\nu}\big{)}.$
By (4.18) and the condition $F_{-}^{*}(\nu_{-})=0$,
[…]
By (5.7) and (5.8),
$L_{\nu}\big{(}e^{\psi\circ F}|F_{\bar{z}}|^{2}\big{)}$
$=$ $(1-|\mu_{F}|^{2})^{-1}\Big{(}\Phi\big{(}2\nu-2\bar{\nu}\mu_{F}^{2}-\nu+\nu|\mu_{F}|^{2}\big{)}+\overline{\Phi}\big{(}2\bar{\nu}-2{\nu}\bar{\mu}_{F}^{2}-\bar{\nu}+\bar{\nu}|\mu_{F}|^{2}\big{)}\Big{)}$
$=$ $e^{\phi}(1-|\mu_{F}|^{2})^{-1}\Big{(}\nu\bar{\mu}_{F}+\nu\bar{\mu}_{F}|\mu_{F}|^{2}-2\bar{\nu}{\mu}_{F}|\mu_{F}|^{2}+\bar{\nu}{\mu}_{F}+\bar{\nu}{\mu}_{F}|\mu_{F}|^{2}-2{\nu}\bar{\mu}_{F}|\mu_{F}|^{2}\Big{)}$
$=$ $e^{\phi}(1-|\mu_{F}|^{2})^{-1}\Big{(}\nu\bar{\mu}_{F}-\bar{\nu}{\mu}_{F}|\mu_{F}|^{2}+\bar{\nu}{\mu}_{F}-{\nu}\bar{\mu}_{F}|\mu_{F}|^{2}\Big{)}$
$=$ $e^{\phi}\Big{(}\nu\bar{\mu}_{F}+\bar{\nu}{\mu}_{F}\Big{)}.$
Hence,
$\frac{1}{2}\big{(}L_{\nu}-iL_{i\nu}\big{)}\big{(}e^{\psi\circ F}|F_{\bar{z}}|^{2}\big{)}=\frac{1}{2}e^{\phi}\Big{(}\nu\bar{\mu}_{F}+\bar{\nu}{\mu}_{F}+\nu\bar{\mu}_{F}-\bar{\nu}{\mu}_{F}\Big{)}=e^{\phi}\nu\bar{\mu}_{F}.$
Proposition 5.4.
Under the condition $F_{+}^{*}(\nu_{+})=0$ or $F_{-}^{*}(\nu_{-})=0$,
(5.9) $\frac{1}{2}\big{(}L_{\mu}+iL_{i\mu}\big{)}\big{(}e^{\phi}\nu\bar{\mu}_{F}\big{)}=e^{\phi}(1+|\mu_{F}|^{2})\nu\bar{\mu}.$
Proof.
We will prove only for the case $F_{-}^{*}(\nu_{-})=0$ since the proof for the other case is the same. By (4.10) and $L_{\mu}\nu=0$ (see (2.3) of [20]),
$L_{\mu}\big{(}e^{\phi}\nu\bar{\mu}_{F}\big{)}=L_{\mu}\big{(}\Phi\nu\big{)}$
$=$ $\,e^{\phi}\big{(}F^{*}({{\mu_{h}}})\bar{\mu}_{F}^{2}-{\bar{\mu}}|\mu_{F}|^{2}\big{)}+e^{\phi}\big{(}F^{*}(\bar{{\mu}}_{h})-\bar{{\mu}}\big{)}\,\nu$
$=$ $\,e^{\phi}(1-|\mu_{F}|^{2})^{-1}\Big{(}2\mu\bar{\mu}_{F}^{2}-2\bar{\mu}|\mu_{F}|^{4}-\bar{\mu}|\mu_{F}|^{2}+\bar{\mu}|\mu_{F}|^{4}\Big{)}\nu$
$+e^{\phi}(1-|\mu_{F}|^{2})^{-1}\Big{(}2\bar{\mu}-2{\mu}\bar{\mu}_{F}^{2}-\bar{\mu}+\bar{\mu}|\mu_{F}|^{2}\Big{)}\nu$
$=$ $e^{\phi}(1-|\mu_{F}|^{2})^{-1}\bar{\mu}(1-|\mu_{F}|^{4})\nu=e^{\phi}(1+|\mu_{F}|^{2})\bar{\mu}\nu.$
Hence,
$\frac{1}{2}\big{(}L_{\mu}+iL_{i\mu}\big{)}\big{(}e^{\phi}\nu\bar{\mu}_{F}\big{)}=e^{\phi}(1+|\mu_{F}|^{2})\nu\bar{\mu}.$
Theorem 5.5.
Over the subspace $T_{0}(1)^{\pm}$ of $T^{*}T_{0}(1)$,
(5.10) ¯$\partial\bar{\partial}\Big{(}2\int_{\Sigma}e^{\psi\circ F}|F_{\bar{z}}|^{2}\,d^{2}z\Big{)}=\mp\sqrt{-1}\,i^{*}_{\pm}\omega_{\mathrm{C}},$
where $i_{\pm}:T_{0}(1)^{\pm}\to T^{*}T_{0}(1)$ denotes the embedding map of $T_{0}(1)^{\pm}$ respectively.
Proof.
We prove only for the case of $T_{0}(1)^{+}$ since the proof of the other case is essentially same.
In order to consider variation of the anti-holomorphic energy of $F_{\pm}$, we have to deal with variation of the anti-holomorphic density $e^{\phi}|\mu_{F}|^{2}\,d^{2}z$ together. For the variation of $d^{2}z=-\sqrt{-1}/2dz\wedge d\bar{z}$ along a family of quasi-conformal map $f^{\epsilon}$, we have
$df^{\epsilon}\wedge d\bar{f^{\epsilon}}=\big{(}|f^{\epsilon}_{z}|^{2}-|f^{\epsilon}_{\bar{z}}|^{2}\big{)}\,dz\wedge d\bar{z}=|f^{\epsilon}_{z}|^{2}\big{(}1-\epsilon^{2}|\mu_{f}|^{2}\big{)}\,dz\wedge d\bar{z}$
where we used the equality $f^{\epsilon}_{\bar{z}}=\epsilon\mu_{f}f^{\epsilon}_{z}$ for the second equality. Hence, the factor $|f^{\epsilon}_{z}|^{2}$ with the variational term of $e^{\phi}|\mu_{F}|^{2}$ precisely gives the Lie derivative of this in Propositions 5.3 and 5.4. Then, we have
(5.11) $\begin{split}\partial_{\nu}\Big{(}2\int_{\Sigma}e^{\psi\circ F}|F_{\bar{z}}|^{2}\,d^{2}z\Big{)}=&2\int_{\Sigma}e^{\phi}\,\nu\bar{\mu}_{F}\,d^{2}z,\\ \bar{\partial}_{\mu}\partial_{\nu}\Big{(}2\int_{\Sigma}e^{\psi\circ F}|F_{\bar{z}}|^{2}\,d^{2}z\Big{)}=&2\int_{\Sigma}e^{\phi}\,(1+|\mu_{F}|^{2})\,\nu\bar{\mu}\,d^{2}z.\end{split}$
By the condition $F_{-}^{*}(\nu_{-})=0$ and (4.24),
(5.12) $\begin{split}\omega_{\mathrm{C}}=&-\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F}|^{2})\,e^{\phi}\,\sum_{k\in\mathbb{Z}}\Big{(}F^{*}_{+}(\nu_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})\Big{)}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F}|^{2})\,e^{\psi\circ F_{+}}|(F_{+})_{z}|^{2}\,\sum_{k\in\mathbb{Z}}\Big{(}F^{*}_{+}(\nu_{+,k})\wedge F^{*}_{+}(\bar{\nu}_{+,k})\Big{)}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\mathrm{Mess}^{*}_{+}\int_{\mathbb{D}}\,e^{\psi}\,\sum_{k\in\mathbb{Z}}\Big{(}\nu_{+,k}\wedge\bar{\nu}_{+,k}\Big{)}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\mathrm{Mess}^{*}_{+}\int_{\mathbb{D}}\,e^{\psi}\,\sum_{k,\ell\in\mathbb{Z}}\Big{(}\nu_{+,k}\wedge\bar{\nu}_{+,\ell}\Big{)}\,d^{2}z.\end{split}$
Here the last equality holds since $\{\nu_{+,k}\}_{k\in\mathbb{Z}}$ is an orthonormal basis. By these equalities and (5.8),
(5.13) $\begin{split}&-\frac{\sqrt{-1}}{2}\mathrm{Mess}^{*}_{+}\int_{\mathbb{D}}\,e^{\psi}\,\Big{(}\nu_{+,k}\wedge\bar{\nu}_{+,\ell}+\nu_{+,\ell}\wedge\bar{\nu}_{+,k}\Big{)}\,d^{2}z\\ =&-\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F}|^{2})\,e^{\phi}\\ &\qquad\cdot\Big{(}2(1-|\mu_{F}|^{2})^{-1}(\nu_{k}-\bar{\nu}_{k}\mu_{F}^{2})\wedge 2(1-|\mu_{F}|^{2})^{-1}(\bar{\nu}_{\ell}-\nu_{\ell}\bar{\mu}_{F}^{2})\Big{)}\,d^{2}z\\ &-\frac{\sqrt{-1}}{2}\int_{\Sigma}(1-|\mu_{F}|^{2})\,e^{\phi}\\ &\qquad\cdot\Big{(}2(1-|\mu_{F}|^{2})^{-1}(\nu_{\ell}-\bar{\nu}_{\ell}\mu_{F}^{2})\wedge 2(1-|\mu_{F}|^{2})^{-1}(\bar{\nu}_{k}-\nu_{k}\bar{\mu}_{F}^{2})\Big{)}\,d^{2}z\\ =&-2{\sqrt{-1}}\int_{\Sigma}(1-|\mu_{F}|^{2})^{-1}\,e^{\phi}\\ &\qquad\qquad\cdot\Big{(}\nu_{k}\wedge\nu_{\ell}+\nu_{\ell}\wedge\bar{\nu}_{k}+\bar{\nu}_{k}\wedge\nu_{\ell}\,|\mu_{F}|^{4}+\bar{\nu}_{\ell}\wedge\nu_{k}\,|\mu_{F}|^{4}\Big{)}\,d^{2}z\\ =&-2\sqrt{-1}\int_{\Sigma}(1+|\mu_{F}|^{2})\,e^{\phi}\,\Big{(}\nu_{k}\wedge\bar{\nu}_{\ell}+\nu_{\ell}\wedge\bar{\nu}_{k}\Big{)}\,d^{2}z.\end{split}$
Combining (5.12) and (5.13),
(5.14) $-\sqrt{-1}\,i^{*}_{+}\omega_{C}=-2\int_{\Sigma}(1+|\mu_{F}|^{2})\,e^{\phi}\,\sum_{k,\ell\in\mathbb{Z}}\Big{(}\nu_{k}\wedge\bar{\nu}_{\ell}\Big{)}\,d^{2}z.$
Hence, by (5.11) and (5.14), we conclude
¯$\partial\bar{\partial}\Big{(}2\int_{\Sigma}e^{\phi}|\mu_{F}|^{2}\,d^{2}z\Big{)}=-\sqrt{-1}\,i^{*}_{+}\omega_{\mathrm{C}}\qquad\text{over}\quad T_{0}(1)^{+}.$
This completes the proof for the case of $T_{0}(1)^{+}$. ∎
By Theorem 4.10 and Theorem 5.5,
Theorem 5.6.
Over the subspace $T_{0}(1)^{\pm}$ of $T^{*}T_{0}(1)$,
¯$\partial\bar{\partial}\Big{(}2\int_{\Sigma}e^{\phi}|\mu_{F}|^{2}\,d^{2}z\Big{)}=\sqrt{-1}\,i^{*}_{\pm}\mathrm{Mess}^{*}_{\pm}\big{(}\omega_{\mathrm{WP}}\big{)}.$
Remark 5.7.
By Remark 4.12 and the proof of Theorem 5.5, the pullback 2-form $\mathrm{Mess}^{*}_{\pm}(\omega_{\mathrm{WP}})$ to $T_{0}(1)^{\pm}$ has the following expression
(5.15) $i^{*}_{\pm}\mathrm{Mess}^{*}_{\pm}(\omega_{\mathrm{WP}})=2{\sqrt{-1}}\int_{\Sigma}(1+|\mu_{F}|^{2})e^{\phi}\,\sum_{k,\ell\in\mathbb{Z}}\,\nu_{k}\wedge\bar{\nu}_{\ell}\ d^{2}z$
where $\{\nu_{k}\}_{k\in\mathbb{Z}}$ is a basis of $T_{[\mu]}T_{0}(1)$.
Appendix A Anti de Sitter space of dimension 3
Let $\mathbb{R}^{2,2}$ denote the pseudo-Euclidean 4-space endowed with linear coordinates $\mathbf{x}=(x_{1},x_{2},x_{3},x_{4})$. Let $q$ denote the quadratic form defined by
(A.1) $q(\mathbf{x})=x^{2}_{1}+x^{2}_{2}-x^{2}_{3}-x^{2}_{4},$
[…]
(A.2) $\mathbb{H}^{2,1}=\{\,\mathbf{x}\in\mathbb{R}^{2,2}\,|\,q(\mathbf{x})=-1\,\}.$
One can check that $\mathbb{H}^{2,1}$ is a smooth connected submanifold of $\mathbb{R}^{2,2}$ of dimension $3$. The tangent space $T_{\mathbf{x}}\mathbb{H}^{2,1}$ can be identified with a subspace of $\mathbb{R}^{2,2}$, which is given by
$\mathbf{x}^{\perp}=\{\mathbf{y}\in\mathbb{R}^{2,2}\,|\,\langle\mathbf{x},\mathbf{y}\rangle=0\}.$
[…]
Let us define the Anti-de Sitter space of dimension 3 by
(A.3) $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}:=\mathbb{H}^{2,1}/\{\pm\mathrm{Id}\}$
where $\mathrm{Id}$ denotes the identity element in $\mathrm{O}(2,2)$. Then $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is a Lorentzian manifold with the induced metric from $\mathbb{H}^{2,1}$ and has the constant curvature $-1$. By the definition, $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is identified with a subset of real projective space $\mathbb{RP}^{3}$, that is
$\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}=\{[\mathbf{x}]\in\mathbb{RP}^{3}\,|\,q(\mathbf{x})<0\}.$
The boundary of $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ in $\mathbb{RP}^{3}$ is the projectivization of the set of lightlike vectors in $\mathbb{R}^{2,2}$,
(A.4) $\partial\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}=\{[\mathbf{x}]\in\mathbb{RP}^{3}\,|\,q(\mathbf{x})=0\}.$
Let $\mathrm{M}(2,\mathbb{R})$ denote the vector space of $2\times 2$ matrices with real entries. There is an isometric identification between $(\mathrm{M}(2,\mathbb{R}),-\mathrm{det})$ and $(\mathbb{R}^{2,2},q)$. Under this isomorphism, $\mathbb{H}^{2,1}$ is identified with the Lie group $\mathrm{SL}(2,\mathbb{R})$. The Lie group $\mathrm{SL}(2,\mathbb{R})$ has the Killing form $\kappa$, that is a bi-invariant bilinear form on the Lie algebra $\mathfrak{sl}(2,\mathbb{R})$ of the signature $(2,1)$. The bilinear form $\kappa$ induces a Lorentzian metric on $\mathrm{SL}(2,\mathbb{R})$, which we denote by $g_{\kappa}$. Recalling (A.3), one can check the Anti-de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is the same as $\mathrm{PSL}(2,\mathbb{R})$ equipped with $\frac{1}{8}g_{\kappa}$ as a Lorentzian manifold.
Let us notice that $\mathrm{SL}(2,\mathbb{R})\times\mathrm{SL}(2,\mathbb{R})$ acts on $\mathrm{M}(2,\mathbb{R})$ by
$(\alpha,\beta)\cdot\gamma=\alpha\circ\gamma\circ\beta^{-1}\qquad\text{for}\quad\gamma\in\mathrm{M}(2,\mathbb{R}),$
and this action preserves the quadratic form $-\mathrm{det}\cong q$. Hence, we have
$\mathrm{Isom}_{0}(\mathbb{H}^{2,1})\cong\mathrm{SO}_{0}(\mathrm{M}(2,\mathbb{R}),q)\cong\large(\mathrm{SL}(2,\mathbb{R})\times\mathrm{SL}(2,\mathbb{R})\large)/K,$
where $K:=\{(\mathrm{Id},\mathrm{Id}),(-\mathrm{Id},-\mathrm{Id})\}$, so that
$\mathrm{Isom}_{0}(\mathbb{A}\mathrm{d}\mathbb{S}^{2,1})\cong\mathrm{PSL}(2,\mathbb{R})\times\mathrm{PSL}(2,\mathbb{R}).$
By (A.4), for the projectivization of $\mathrm{M}(2,\mathbb{R})$ denoted by $\mathrm{P}(\mathrm{M}(2,\mathbb{R}))$,
(A.5) $\partial\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}=\{[X]\in\mathrm{P}(\mathrm{M}(2,\mathbb{R}))\,|\,\mathrm{rank}(X)=1\}.$
We have the following homeomorphism
(A.6) $\partial\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}\,\to\,\mathbb{R}\mathbb{P}^{1}\times\mathbb{R}\mathbb{P}^{1}$
defined by
(A.7) $[X]\mapsto(\mathrm{Im}(X),\mathrm{Ker}(X)).$
A timelike geodesic in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ has the following form
$L_{p,q}=\{\gamma\in\mathrm{PSL}(2,\mathbb{R})\,|\,\gamma(q)=p\,\}$
for some points $p,q\in\mathbb{D}$.
For a spacelike conformal embedding $\sigma:\mathbb{D}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, let $\Sigma$ denote the image $\sigma(\mathbb{D})\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. The associated Gauss map
(A.8) $G:\Sigma\to\mathbb{D}\times\mathbb{D}$
is defined by $G(x)=(p,q)$ where $L_{p,q}$ is the timelike geodesic orthogonal to the tangent space of $\Sigma$ at $x$.